### v2

# Deep Submillimeter and Radio Observations in the SSA22 Field. III. Multiwavelength Identifications and Properties of the 850 -selected Submillimeter Galaxies

Abstract
We present a multiwavelength analysis of 850 -selected submillimeter galaxies (SMGs; deblended S 1 mJy) in the SSA22 field, using JCMT/SCUBA-2 data reaching 0.80 mJy beam-1. We identify 248 SMG candidates for 192 SCUBA-2 sources, with a multiplicity rate of 26% that increases with flux. After quality cuts based on SED fitting reliability, our final sample contains 221 SMGs associated with 186 SCUBA-2 sources. The SSA22 SMGs have a median infrared luminosity of (2.25 0.25) 1012 L⊙, with 63% ( 8%) of the sample classified as ULIRGs (HLIRGs). The median redshift is (higher, , for optically faint galaxies). SMGs trace large-scale structure, showing a significant overdensity at . The comoving SMG number density rises by a factor of at , plateauing at cMpc-3 over 1-3 (including the overdensity). The sample has median stellar mass M⊙, SFR M⊙ yr-1, dust mass M⊙, and mag. We detect a clear “downsizing” signature: , massive SMGs quench, while lower-mass systems dominate later star formation. The dust-to-stellar mass ratio, median , evolves significantly below . Although remains stable over , combined literature data suggest a sharp drop at , indicating a transition to less obscured star formation likely due to low metal enrichment in the early universe.
I Introduction
Massive dusty star-forming galaxies represent crucial probes for addressing fundamental questions about our universe: when and how did most stars and galaxies form and evolve? How did cosmic metals enrich? What was the assemble history and dynamic evolution of the galaxy? How did the large-scale structure form and interact with galaxies? Following cosmic reionization, dust-obscured star formation gradually dominated the cosmic star formation rate density, contributing approximately 50-80% of the total (Bourne et al., 2017; Dunlop et al., 2017; zavala2021). However, such activity—hidden within dense dust clouds—often evades detection at ultraviolet and optical wavelengths (wang2019; Hyun et al., 2023), making direct observation of dust heated by intense star formation a key method for its study (Hughes et al., 1998).
Ultraviolet photons from massive stars are absorbed by dense dust and re-emitted in the far-infrared (peak wavelength 100 ; Lim et al., 2020b), resembling modified blackbody radiation (Dudzevičiūtė et al., 2020b). Submillimeter observations lie on the Rayleigh–Jeans tail of this emission. As redshift increases, the observed band shifts toward the peak, counteracting cosmological dimming via the negative -correction (Blain, 2002; Casey et al., 2014). Longer wavelengths enhance this effect, leading to higher typical redshifts for SMGs selected at longer wavelengths (Chapin et al., 2009; Chen et al., 2013a; Roseboom et al., 2013; Casey et al., 2014; Zavala et al., 2014; Wang et al., 2017). This work focuses primarily on 850 -selected SMGs. Observational constraints and atmospheric windows have historically confined submillimeter surveys to around 850 . At this wavelength, the negative -correction nearly balances cosmic dimming, keeping the observed flux density relatively constant over for galaxies of similar luminosity (Blain, 2002). Thus, submillimeter surveys allow the construction of samples of dust-obscured star-forming galaxies across a wide cosmic timeframe (Dudzevičiūtė et al., 2020b).
The study of SMGs has advanced considerably over the past three decades. Initial detections with JCMT/SCUBA in the late 1990s (Smail et al., 1997; Barger et al., 1998; Hughes et al., 1998; Ivison et al., 1998; Lilly et al., 1999) revealed a population of galaxies hidden by dust. Large-area surveys like the SCUBA-2 Cosmology Legacy Survey (S2CLS; Geach et al. 2017) and the SCUBA-2 Large eXtragalactic Survey (S2LXS; Garratt et al. 2023), covering 5-10 deg2, along with deep fields such as STUDIES in COSMOS field (Wang et al., 2017; Gao et al., 2024), have expanded our understanding. Single-dish submillimeter/millimeter surveys are limited by confusion noise and sensitivity, typically resolving only 10–40% of the cosmic infrared background (CIB) (Wang et al., 2017; Gao et al., 2024). Gravitational lensing has enabled the study of faint SMGs (Hsu et al., 2016), with targeted lensing-field surveys resolving 40–60% of the CIB—and in some cases, up to 80% (Chen et al., 2013a, b). We anticipate that next-generation submillimeter facilities, such as the Xue-shan-mu-chang 15-meter SubMillimeter Telescope (XSMT-15m, XSMT Project Collaboration Group et al., 2025), will improve constraints on both the cosmic infrared background and the cosmic star formation rate density (CSFRD).
Facilities like Spitzer, Herschel, IRAM 30-m, ASTE, APEX, and SPT have greatly contributed, though their coarse angular resolution (Hodge and da Cunha, 2020) has been a limitation. Interferometers like SMA, NOEMA, and especially ALMA have revolutionized the field with high resolution and sensitivity, enabling detailed studies of individual high-redshift SMGs, including their gas dynamics (Hodge and da Cunha, 2020). Umehata et al. (2017, 2018); Franco et al. (2018); Hatsukade et al. (2018); Gómez-Guijarro et al. (2022) mapped GOODS-S and SSA22 in a wide region utilizing ALMA, while the Ex-MORA survey has assembled the largest homogeneous ALMA continuum dataset to date in the COSMOS field (Casey et al., 2021; Long et al., 2024). Typical ALMA surveys achieve CIB resolution fractions of 40–50% (Chen et al., 2022), and—again leveraging gravitational lensing—the ALCS survey successfully resolved 80% of the CIB (Fujimoto et al., 2024). The UDS and COSMOS fields host the largest samples of ALMA follow-up observations (Stach et al., 2019; Dudzevičiūtė et al., 2020b; simpson2020; Liu et al., 2019; Adscheid et al., 2024). Collectively, these efforts have expanded SMG studies in depth, breadth, and methodological sophistication.
These observations have revealed that SMGs are a heterogeneous population, exhibiting significant scatter relative to the star-forming main sequence (Shim et al., 2022), yet they are consistently massive, dust-obscured star-forming galaxies. 850 -selected SMGs typically possess stellar masses of M⊙, with a median near M⊙, and reside in massive dark matter halos ( M⊙; Lim et al., 2020b; Stach et al., 2021). Their star formation rates (SFRs) range from tens to thousands of solar masses per year. Sustaining such an active process requires substantial gas reservoirs, with typical SMGs hosting gas masses of M⊙ and gas fractions 0.5-0.9 (miettinen2017; birkin2021; liao2024), fueled either by major mergers or environmental accretion. Their high redshift, dust obscuration, and intense star formation result in extremely red colors and high infrared luminosities L⊙, with cold dust temperatures of 20-40 K (swinbank2014; Dudzevičiūtė et al., 2020b). The fraction of SMGs hosting active galactic nuclei (AGNs) has been a subject of debate in earlier studies. Initial work suggested AGN fractions as high as 20–40% (Alexander et al., 2005a; Pope et al., 2008; Laird et al., 2010; Georgantopoulos et al., 2010; Johnson et al., 2013; Wang et al., 2013). However, studies in the ALMA era have revealed that these earlier estimates were likely inflated. More recent analyses indicate that the AGN fraction among SMGs is below 20%, typically in the range of 15-20% (Wang et al., 2013; An et al., 2019; Stach et al., 2019; uematsu2025). Moreover, brighter and more massive SMGs appear more likely to host AGNs (Wang et al., 2013; Cowie et al., 2018; Lim et al., 2020b). SMGs residing in high-redshift large-scale structures may also exhibit elevated AGN fractions (Umehata et al., 2015). Additionally, SMGs hosting AGNs show a higher incidence of major mergers (uematsu2025), suggesting that major mergers play a key role in triggering black hole activity in these systems. In some cases, SMGs are also identified as optically-selected quasars, where concurrent black hole growth and star formation drive rapid galaxy evolution (Alexander et al., 2005a, b). These characteristics strongly suggest that SMGs are progenitors of present-day massive elliptical galaxies, which hosting supermassive black holes (SMBHs) at their centers (Genzel et al., 2003; simpson2017; Lim et al., 2020a).
Observations suggest that SMGs are often associated with overdense structures at high redshift ( 1.5 ) (Menéndez-Delmestre et al., 2013; Dannerbauer et al., 2014; Arrigoni Battaia et al., 2018; Smail, 2024; Zhang et al., 2024), and broadly consistent with the general tendency for massive galaxies to reside in denser environments (Dressler, 1980). Umehata et al. (2015) first demonstrated—free from source confusion—that SMGs anchor galaxy overdensities within cosmic web filaments, and Umehata et al. (2019) later revealed the co-evolution of galaxies, gas, and black holes within these cosmic structures. Thus, SMGs have been considered potential tracers of overdense regions (Tamura et al., 2009; Umehata et al., 2019), though their utility may decline below due to “downsizing” and decreasing baryon conversion efficiency (Chapman et al., 2009; Miller et al., 2015, 2018), and they may incompletely trace the most massive overdensities (Miller et al., 2015).
The formation mechanisms of SMGs are debated. Early views, influenced by their resemblance to scaled-up local ULIRGs, favored major gas-rich mergers (swinbank2014; Hyun et al., 2023), supported by morphological, kinematic, and simulation studies (Toft et al., 2014). Alternatively, some argue that mergers may be secondary (Daddi et al., 2007). Many SMGs lie on the high-mass end of the star-forming main sequence (cunha2015; Dunlop et al., 2017; Zavala et al., 2018; Lim et al., 2020b). Recent ALMA and JWST studies find that of SMGs possess undisturbed extended disks (Gillman et al., 2023, 2024; Le Bail et al., 2024; Chan et al., 2025; Hodge et al., 2025; Umehata et al., 2025a; Zhang et al., 2025), suggesting that internal instabilities or minor perturbations can trigger intense star formation and dust production in isolated disks (Gillman et al., 2024). Numerical models also support sustained starbursts in gas-rich disks (swinbank2014).
The S2CLS is a major JCMT/SCUBA-2 program that mapped seven extragalactic deep fields, covering a total area 5 deg2 and a depth of mJy (Geach et al., 2017). SCUBA-2 sources in fields such as COSMOS (An et al., 2019; simpson2019; simpson2020), UDS (Stach et al., 2018, 2019; Dudzevičiūtė et al., 2020b), EGS (Zavala et al., 2018), and NEP (Shim et al., 2022) have been extensively analyzed. As one of the S2CLS target fields, SSA22 field was well observed; however, to better exploit its potential for studying SMGs, we obtained deeper SCUBA-2 integrations toward SSA22 and complemented them with coordinated VLA 3 GHz observations (Ao et al., 2017; Zeng et al., 2024). The SSA22 field has been extensively studied over the past two decades across various wavelengths and cosmic structures (C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 113; Y. Matsuda, T. Yamada, T. Hayashino, H. Tamura, R. Yamauchi, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2005); B. D. Lehmer, D. M. Alexander, J. E. Geach, I. Smail, A. Basu-Zych, F. E. Bauer, S. C. Chapman, Y. Matsuda, C. A. Scharf, M. Volonteri, and T. Yamada (2009b); B. D. Lehmer, D. M. Alexander, S. C. Chapman, I. Smail, F. E. Bauer, W. N. Brandt, J. E. Geach, Y. Matsuda, J. R. Mullaney, and A. M. Swinbank (2009a); T. M. A. Webb, T. Yamada, J.-S. Huang, M. L. N. Ashby, Y. Matsuda, E. Egami, M. Gonzalez, and T. Hayashimo (2009); M. Kubo, Y. K. Uchimoto, T. Yamada, M. Kajisawa, T. Ichikawa, Y. Matsuda, M. Akiyama, T. Hayashino, M. Konishi, T. Nishimura, K. Omata, R. Suzuki, I. Tanaka, T. Yoshikawa, D. M. Alexander, G. G. Fazio, J.-S. Huang, and B. D. Lehmer (2013); M. Kubo, T. Yamada, T. Ichikawa, M. Kajisawa, Y. Matsuda, I. Tanaka, and H. Umehata (2015); H. Umehata, Y. Tamura, K. Kohno, B. Hatsukade, K. S. Scott, M. Kubo, T. Yamada, R. J. Ivison, R. Cybulski, I. Aretxaga, J. Austermann, D. H. Hughes, H. Ezawa, T. Hayashino, S. Ikarashi, D. Iono, R. Kawabe, Y. Matsuda, H. Matsuo, K. Nakanishi, T. Oshima, T. Perera, T. Takata, G. W. Wilson, and M. S. Yun (2014); H. Umehata, M. Fumagalli, I. Smail, Y. Matsuda, A. M. Swinbank, S. Cantalupo, C. Sykes, R. J. Ivison, C. C. Steidel, A. E. Shapley, J. Vernet, T. Yamada, Y. Tamura, M. Kubo, K. Nakanishi, M. Kajisawa, B. Hatsukade, and K. Kohno (2019); H. Umehata, M. Kubo, I. Smail, B. D. Lehmer, E. B. Monson, K. Nakanishi, and Y. Matsuda (2025a); Y. Kato, Y. Matsuda, I. Smail, A. M. Swinbank, B. Hatsukade, H. Umehata, I. Tanaka, T. Saito, D. Iono, Y. Tamura, K. Kohno, D. K. Erb, B. D. Lehmer, J. E. Geach, C. C. Steidel, D. M. Alexander, T. Yamada, and T. Hayashino (2016); Y. Ao, Y. Matsuda, C. Henkel, D. Iono, D. M. Alexander, S. C. Chapman, J. Geach, B. Hatsukade, M. Hayes, N. K. Hine, Y. Kato, R. Kawabe, K. Kohno, M. Kubo, M. Lehnert, M. Malkan, K. M. Menten, T. Nagao, R. P. Norris, M. Ouchi, T. Saito, Y. Tamura, Y. Taniguchi, H. Umehata, and A. Weiss (2017); O. R. Cooper, C. M. Casey, J. A. Zavala, J. B. Champagne, E. da Cunha, A. S. Long, J. S. Spilker, and J. Staguhn (2022); X. Zeng, Y. Ao, and Y. Zhang (2024); S. Huang, H. Umehata, I. Smail, K. Nakanishi, B. Hatsukade, M. Kubo, Y. Tamura, T. Saito, and S. Ikarashi (2025)), amassing rich multiwavelength data from X-ray to radio. It offers a unique laboratory for probing both early galaxy evolution and the assembly of cosmic web structures.
In this paper, we combine 850 and 3 GHz data with extensive multi-wavelength ancillary datasets to identify SMG counterparts, perform spectral energy distribution (SED) fitting, and investigate their physical properties. A key challenge is the coarse resolution of single-dish submillimeter telescopes like SCUBA-2, requiring counterpart identification at higher-resolution. Although bands with positive -corrections and varying depths can miss higher-redshift or warmer SMGs (swinbank2014), they successfully identify the majority within the expected redshift range. Radio and 24 identifications are reliable and widely used (Chen et al., 2016), although these methods still miss a fraction of real SMGs (Hodge et al., 2013; An et al., 2018). The ADF22 survey (Umehata et al., 2017, 2018) provides additional counterparts for our sample. To enhance the completeness and maximize the potential of multiwavelength data, IRAC 8 sources have also been widely utilized (Ashby et al., 2006; Zavala et al., 2018). Numerous studies employ various optical/near-infrared (NIR) wavelengths (Smail et al., 1999, 2002; Frayer et al., 2004) to identify high-redshift extremely red galaxies (e.g., HIEROs: Wang et al. (2012); KIEROs: Wang et al. (2016); OIRTC: Chen et al. (2016)).
Far-infrared (FIR) and millimeter data also suffer from source confusion due to large beam sizes. A critical step in deblending is constructing a prior source catalog that is both complete and optimally dense (Liu et al., 2018). We use the robust “super-deblending” technique (Liu et al., 2018; Jin et al., 2018), which employs SED-based flux predictions to create appropriate prior lists for each band and estimates deblended flux uncertainties via Monte Carlo simulations.
We perform SED modeling using CIGALE, a modular code that combines physical components. By exploring a parameter grid and enforcing energy balance, whereby dust reprocesses ultra-violet (UV) and optical energy from stars and AGN into the far-infrared, it constructs possible SEDs. This full-SED fitting method more effectively breaks degeneracies in galaxy properties compared to using UV/optical or FIR data solely. Crucially, for SMGs lacking spectroscopic redshifts or detectable optical emission—sources often excluded in traditional analyses—full SED fitting provides robust constraints on physical properties and enables reliable photometric redshift estimation (Dudzevičiūtė et al., 2020b).
This paper is structured as follows. In Section II, we describe the submillimeter, VLA, optical/near-infrared, and ancillary data in the SSA22 field, along with data reduction and the super-deblending procedure. Section III details our counterpart identification methods, results, and SMG multiplicity. In Section IV, we derive spectroscopic redshifts from a representative catalog (Mawatari et al., 2023), photometric redshifts using EAZY, and simultaneously obtain SED-based redshifts and physical properties via CIGALE fitting. Finally, Section V presents the statistical properties of SMGs in the SSA22 deep field. Specifically, in Section V.1 we present the redshift distribution of our 850 -selected SMG sample. Section V.2 details the star formation rates, stellar masses, and mass-to-light ratios of these SMGs. Section V.3 shows the distribution of infrared luminosities, while Section V.4 reports on dust masses and dust fractions. In Section V.5, we present the distribution of dust attenuation () and examine its relationship with optical colors.
We adopt the cosmological parameters: H0 = 70 km-1 s-1 Mpc-1, = 0.70, = 0.30. The article employs the Chabrier (2003) initial mass function (IMF).
II Data and Data Reduction
In this section, we will introduce the multiwavelength data we have collected, spanning from optical to radio wavelengths. The majority of the data covers the main region of SSA22, specifically SSA22-Sb1 (Yamada et al., 2012). The coverage areas for each wavelength are illustrated in Figure 1, and details regarding the bands, resolutions, and observational depths for each dataset are provided in Table 1.
II.1 SCUBA-2 850
Here, we provide a brief overview of the JCMT/SCUBA-2 850 m observations. Detailed information on observations, data processing, and source number counts can be found in Zeng et al. (2024).
The SSA22 field was part of the S2CLS (Geach et al., 2017). We conducted additional exposure in the SSA22a region (Ao et al., 2017), which is the core zone of the SSA22 protocluster (Steidel et al., 1998). The total observations expended times of 91 hr and covered an area of 0.34 deg2, reaching a noise level of 0.79 mJy beam-1 in the overlapped area, making it the deepest map at 850 in the SSA22 field (see Figure 2).
We utilize a stacked point-spread function (PSF) and a “top-down” iterative source extraction algorithm. Jackknife maps are constructed, and Monte Carlo simulations are employed to estimate the completeness and false detection rates. Compared to the S2CLS SSA22 map, our deeper SCUBA-2 map reveals 212 new sources, of which 145 lie in the central region of the field (where 1.2 mJy) and 122 have flux densities greater than 4 mJy. While most of these newly detected sources are at low signal-to-noise ratios (SNR ), we still recover 93 sources at significance. Accounting for the false detection rate (Zeng et al., 2024), we estimate that approximately 170 of these new detections are authentic submillimeter sources.
Finally, we construct a catalog comprising 390 SCUBA-2 sources with significance 3.5111Although sources below 4 carry an elevated risk of being spurious, statistical considerations suggest that 60-80% of them are likely real. and flux densities 2 mJy.
II.2 JVLA 3 GHz
Around our SCUBA-2 observations, we conducted JVLA 3 GHz radio band observations with B-configuration at S-band (2–4 GHz) in 2015 and 2016 (Ao et al., 2017). The observations combined three pointings, and the primary beam of a single observation is about 15. The final mosaic of images achieved an sensitivity of 1.5 Jy beam-1 before primary beam correction, with an angular resolution of 23 20.
We used the Python Blob Detection and Source Finder software (PyBDSF v1.10.3; Mohan and Rafferty, 2015) to detect and extract the radio sources from the JVLA 3 GHz map. Using the flux and root-mean-square (rms) maps after primary beam correction, we applied a sigma clipping threshold to extract the highly significant sources. Pixels above 3 times the background rms were identified as an island, and multiple Gaussians were fitted to the continuous pixels above detection threshold of 5 within each island (Figure 3 panel (a) and (b)). If all pixel values between two peak pixels exceed 3, and the connecting line length is less than half the sum of their FWHM values, these peak pixels are grouped as a single source and modeled with a Gaussian profile. The total flux of each source was estimated by integrating all pixel values within the Gaussian, and the positions of the sources were computed using moment analysis (Jiménez-Andrade et al., 2019; Van Der Vlugt et al., 2021).
Considering that SMGs are generally at high redshifts and have relatively compact starburst regions (Simpson et al., 2015), we did not consider extended emissions. The parameter flag_maxsize_fwhm was set to default value of 0.5, meaning any fitted Gaussian profile with an FWHM exceeding half the island size would be flagged222https://pybdsf.readthedocs.io/en/latest/process_image.html (Figure 3 panel (c)).
Additionally, we observed artifacts around some bright sources caused by projection effects, contamination from nearby faint sources, or noise. These appeared as unnatural extended shadows around high-significance detections.
After visual inspection of the modeled images and flagged sources (Figure 3 panel (b),(c)), we excluded both the flagged sources and two spurious detections, resulting in a total of 1989 radio sources with a significance , which were later used to identify deblended SMGs.
II.3 Optical/Near-infrared
The optical and near-infrared (opt/NIR) multi-band images are centered on the Lyman- emitters (LAEs) overdensity peak region (C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 113; T. Hayashino, Y. Matsuda, H. Tamura, R. Yamauchi, T. Yamada, M. Ajiki, S. S. Fujita, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2004); Y. Matsuda, T. Yamada, T. Hayashino, H. Tamura, R. Yamauchi, T. Murayama, T. Nagao, K. Ohta, S. Okamura, M. Ouchi, K. Shimasaku, Y. Shioya, and Y. Taniguchi (2005); T. Yamada, Y. Nakamura, Y. Matsuda, T. Hayashino, R. Yamauchi, N. Morimoto, K. Kousai, and M. Umemura (2012)) (Figure 4). For a detailed description and data processing of the dataset, please refer to Mawatari et al. (2023).
The dataset includes the following multi-band data: the CFHT/MegaPrime (Boulade et al., 2003) u*-band data; the Subaru/Suprime-Cam (Miyazaki et al., 2002) B, V, R, i′, z′ band images (Hayashino et al., 2004; Matsuda et al., 2004); the Subaru/HSC (Miyazaki et al., 2012, 2018) g, r, i, z, Y band images collected from the Subaru Strategic Program (HSC-SSP; Aihara et al., 2018); and the Subaru/MOIRCS (Ichikawa et al., 2006) J, H, Ks band data (Uchimoto et al., 2008, 2012). Additionally, the dataset includes the UKIRT/WFCAM (Casali et al., 2007) J, K band data from the deep extragalactic survey, which was part of the UKIRT Infrared Deep Sky Survey (UKIDSS/DXS; Lawrence et al., 2007). The two groups of near-infrared bands are similar, but the footprints of WFCAM are more comprehensive, and the MOIRCS bands are deeper.
We used SExtractor (Bertin and Arnouts, 1996) for source extraction and photometry on the optical and near-infrared images. Only a brief description is provided here; more detailed procedures are given in the Appendix A. The image is divided into a grid, and the median and standard deviation of the pixel value distribution are computed for each cell to estimate the local background and RMS. After smoothing and applying spline interpolation, a continuous model of the sky background is constructed and subtracted from the original image, while the resulting RMS map is used for source detection via sigma clipping. Signals are identified as potential sources only if they exceed twice the background noise level and contain at least five connected pixels (Hayashino et al., 2004; Umehata et al., 2014). SExtractor can also deblends overlapping sources by estimating the contribution from neighboring objects—a step particularly useful for faint sources contaminated by bright foreground objects. We adopted the Kron aperture (MAG_AUTO) for automatic photometry to measure source fluxes (Uchimoto et al., 2012), which is regarded as one of the most accurate, adaptive, and robust flux estimation methods and is widely used in extragalactic surveys (Bertin and Arnouts, 1996). SExtractor further refines photometry by estimating the mean local background residual around each signal based on neighboring pixels. The final source catalog is restricted to detections at significance. The diagnostic images for each procedural step are presented in Figure 5.
The photometry was corrected for Galactic extinction, with = 3.1 applied. The color excesses at each position were obtained using the python package dustmap333https://dustmaps.readthedocs.io/en/latest/index.html (Green, 2018), which provides 2D reddening maps of interstellar dust based on the “SFD model” (Schlegel et al., 1998), with a factor of 0.86 also applied to the reddening values (Schlafly and Finkbeiner, 2011). The extinction curves derived by Fitzpatrick (1999) and improved by Indebetouw et al. (2005) in the infrared were selected to derive the extinction at each band (Rodrigo et al., 2024)444http://svo2.cab.inta-csic.es/svo/theory/fps3/index.php?mode=browse.
In Section III.1.2, We identified galaxies with extreme red colors: 2 as potential counterparts to SMGs. The color measurement was conducted on PSF-matched images that were smoothed to the largest PSF within the dataset, all having a Full Width at Half Maximum (FWHM) of approximately 1.1. Given the reduced impact of dust extinction in the near-infrared bands, we employed a 2.2 aperture for photometry on the -band images (WFCAM/ and MOIRCS/ bands), while processing the Scam band using the dual mode of SExtractor with identical configuration. To ensure the reliability of this approach, we focused on sources detected at 4 significance in the band, simultaneously reducing the spatial density of potential candidates.
Finally, we estimate the depth of the opt/NIR imaging data. However, on small scales, pixel-to-pixel correlations may arise from either source confusion or systematic biases introduced during data reduction. Similarly, on large scale, pixel correlations can result from imperfect background subtraction, contamination by extended sources, undetected faint sources, or systematic errors such as flat-fielding inaccuracies (Straatman et al., 2016). Hence, we perform aperture photometry 10000 times using a 2 radius on background-subtracted and detected sources masked images in each band, generated by SExtractor. The photometry values represent noise distribution, from which we calculate their standard deviation and convert it to magnitudes as the 1 depth of the images, referred to as the 1 limiting magnitudes (Matsuda et al., 2004; An et al., 2018; Weaver et al., 2022). Based on this, we estimate the corresponding 3 depth of the images for each band, which are summarized in Table 1, along with additional multi-band data depths, resolutions.
| Instrument | Band | Angular resolution | 5 limiting mag. | Reference |
| (arcsec) | (AB mag) | |||
| CFHT/MagePrime | u∗ | 1.04 | 26.8 | Kousai (2011) |
| Subaru/Suprime-Cam | B | 0.78 | 27.0 | Nakamura et al. (2011) |
| V | 0.82 | 27.2 | Nakamura et al. (2011) | |
| R | 1.08 | 27.1 | Hayashino et al. (2004) | |
| i′ | 0.76 | 26.6 | Nakamura et al. (2011) | |
| z′ | 0.76 | 26.1 | Nakamura et al. (2011) | |
| Subaru/HSC | g | 0.63 | 26.2 | Aihara et al. (2018) |
| r | 0.74 | 25.6 | Aihara et al. (2018) | |
| i | 0.63 | 25.6 | Aihara et al. (2018) | |
| z | 0.84 | 24.7 | Aihara et al. (2018) | |
| Y | 0.56 | 24.2 | Aihara et al. (2018) | |
| Subaru/MOIRCS | J | 0.68 | 24.6 | Uchimoto et al. (2012) |
| H | 0.68 | 23.9 | Uchimoto et al. (2012) | |
| Ks | 0.55 | 23.9 | Uchimoto et al. (2012) | |
| UKIRT/WFCAM | J | 0.91 | 23.8 | Lawrence et al. (2007) |
| K | 0.86 | 23.2 | Lawrence et al. (2007) | |
| Spitzer/IRAC | 3.6 | 1.7 | 26.00/22.49 ccThe mean 3 noise levels of SEPI and SpIES. | SSC And IRSA (2020); Timlin et al. (2016) |
| 4.5 | 1.7 | 25.24/22.56 ccThe mean 3 noise levels of SEPI and SpIES. | SSC And IRSA (2020); Timlin et al. (2016) | |
| 5.8 | 1.9 | 23.68 ddThe mean 3 noise level of SEPI. | SSC And IRSA (2020) | |
| 8.0 | 2.0 | 22.08 ddThe mean 3 noise level of SEPI. | SSC And IRSA (2020) | |
| Spitzer/MIPS | 24 | 5.9 | 45.6 Jy beam-1 ddThe mean 3 noise level of SEPI. | SSC And IRSA (2020) |
| Herschel/SPIRE | 250 | 17.9 | 2.6 mJy beam-1 eeThe median 1 depth of HHLI data. | HSA (2020) |
| 350 | 24.2 | 2.2 mJy beam-1 eeThe median 1 depth of HHLI data. | HSA (2020) | |
| 500 | 35.4 | 2.7 mJy beam-1 eeThe median 1 depth of HHLI data. | HSA (2020) | |
| JCMT/SCUBA-2 | 850 | 13.9 | 0.8–2.0 mJy beam-1 ffThe 1 depth range of image. | Zeng et al. (2024) |
| AzTEC/ASTE | 1.1 mm | 28 | 0.7–1.3 mJy beam-1 ffThe 1 depth range of image. | Umehata et al. (2014) |
| ALMA | 1.14 mm | 053 052 ggThe resulting synthesized beam size of ALMA band 6. | 75 Jy beam-1 | Umehata et al. (2018) |
| JVLA | 3 GHz | 23 20 hhThe final synthesized beam of VLA S-band. | 1.5 Jy beam-1 iiThe typical sensitivity before primary beam correction. | Ao et al. (2017) |
II.4 Ancillary Data
II.4.1 Mid-infrared
The Spizter data (see also Webb et al., 2009) from the Spitzer Enhanced Imaging Products (SEIP; SSC And IRSA, 2020) and the Spitzer IRAC Equatorial Survey (SpIES; Timlin et al., 2016; Timlin, John D. and Ross, et al, 2020) were utilized.
The SEIP “Super Mosaics” were created by combining multiple observation requests, containing the four channels of IRAC (3.6, 4.5, 5.8, 8 ) and the 24 channel of MIPS. This program produced a photometry source catalog for compact sources and implemented stringent criteria for the selection of extracted sources to ensure high reliability. The mean 3 depths of the four IRAC band enhanced images were approximately 26.00, 25.24, 23.68, 22.08 mag, and the 5 depth for MIPS was approximately 19.20 mag (see Table 1), whereas the SEIP images covered about half of our SCUBA-2 map footprint ( 0.17 and 0.15 deg2 for IRAC and MIPS).
The SpIES is a large-area survey in the equator, fully covering the SSA22 field, thus serve as a supplement to the SEIP. However, the SpIES products only contain the IRAC 3.6 and 4.5 data, and the mapped sensitivities are much shallower than the former. The 3 noise levels for the 3.6 and 4.5 bands are approximately 22.49 and 22.56 AB mag.
We utilized the primary fluxes of the SEIP and SpIES source list, which are the aperture fluxes in a 3.8 diameter of IRAC sources applied aperture correction for point sources, and the PSF fitting fluxes of MIPS 24 sources. To ensure robust flux densities, we excluded sources that could be affected by nearby saturated sources or a nearby extended source and by soft saturation555https://irsa.ipac.caltech.edu/docs/knowledgebase/spitzer_seip.html. Additionally, SNR cuts were applied, with a SNR cut of 3 for IRAC and 5 for MIPS. (FLUX/FLUXERR; Timlin et al., 2016).
II.4.2 Far-infrared and Millimeter
The Herschel data (Kato et al., 2016) from the Herschel High Level Images (HHLI; HSA, 2020), a data subset of the Herschel Science Archive (HSA), included only SPIRE data in the SSA22 field with the median 1 sensitivities were 2.6, 2.2, and 2.7 mJy beam-1 for 250, 350, and 500 , respectively. Due to the large beam of the SPIRE, the three FIR images suffered from severe confusion noise and source blending. The flux densities of SPIRE sources were deblended by the “super deblending” method described in section II.5, and the final 3 deblended sources were used for SED fitting.
Tamura et al. (2009); Umehata et al. (2014) successively conducted the AzTEC/ASTE 1.1 mm survey in the SSA22 field, with a total observation time of 74 hr. The resulting map covered an area of 0.27 deg2, with the rms ranging from 0.7-1.3 mJy beam-1. Considering the coarse resolution of 28 of AzTEC, we also employed the same “super-deblended” program to deblend AzTEC sources to derive 1.1 mm fluxes for identifications for SED fitting.
The ALMA deep field in SSA22 (ADF22 survey; Umehata et al., 2017, 2018) mosaicked a 20 arcmin2 core region at 1.1 mm (band 6), composed of two contiguous observations. The total on-source observation time was approximately 13.1 hr, resulting in a map with a typical sensitivity level of approximately 0.07 mJy beam-1 and resolutions of approximately 0.5 to 1. A total of 35 ALMA SMGs were detected at 5, and the high resolution and precise location of the ALMA observation will be used to identify the positions of the SCUBA-2 sources.
II.5 “Super-deblending”
Accurate FIR to millimeter photometry is essential for measuring star formation rates and dust properties in galaxies, but coarse angular resolution ( 14 - 35) in FIR/mm imaging leads to severe source confusion, especially in deep fields.
We applied the “super-deblending” technique (Liu et al., 2018; Jin et al., 2018) to derive reliable deblended fluxes from 250 to 1.1 mm for SMGs in the SSA22 field. The deblending workflow is presented in Figure 6. We provide a introduction of the “super-deblending” in the Appendix B; for full technical details, please refer to Liu et al. (2018); Jin et al. (2018).
The prior source catalog is constructed from relatively high SNR MIPS 24 and radio sources, and also includes lower-significance ( 2.5) sources at higher redshifts ( 2), resulting in a final catalog of 2683 prior sources.
The deblending is performed individually for each band. A key aspect of the method is the use of SED-fitting predictions to optimize the prior source catalog, thereby enabling reliable flux extraction. To obtain meaningful deblended fluxes, the source density within each FIR/mm beam must be appropriately controlled, ideally kept near or below one source per beam area. For each band, a flux threshold S is defined. Prior sources whose SED-predicted flux (accounting for fitting uncertainties) falls below the threshold considered negligible contributors at that band and are temporarily excluded. A blind extraction is also performed on the residual image to avoid missing any significant FIR emitters. The deblending proceeds sequentially from the shortest to the longest wavelength band. In each step, the deblended fluxes from previously processed bands are used to refine the priors for the current band, which is a crucial feature of the super-deblending method. Finally, Monte Carlo simulations are used to assess potential biases and apply flux corrections, representing a key advantage of the super-deblending approach.
III COUNTERPART IDENTIFICATION
We primarily utilize four methods, namely VLA 3 GHz, /MIPS 24 , /IRAC 8 , and optical/NIR red-selected galaxies (hereafter “/radio”,“”, “”, “” method), to search for potential counterpart galaxies of SCUBA-2 sources.
The radio and identification methods (Section III.1.1) are the most reliable and widely used approaches for pinpointing SMG counterparts (Barger et al., 2000; Ivison et al., 2002; Umehata et al., 2014; Zavala et al., 2018; Lim et al., 2020b), achieving positional accuracies of 50–80% (Hodge et al., 2013; Chen et al., 2016; An et al., 2018). However, their completeness is somewhat lower, typically 40–55% 666Notably, Chen et al. (2016) point out that radio identification alone achieves a higher completeness of 80%.. ADF22 (Umehata et al., 2017, 2018) contributed 5 additional SMG counterparts (two of which were also identified by the method). In subsequent analyses, we designate sources identified through these methods as ‘MAIN’ samples. To fully exploit the potential of multiwavelength data, we also incorporated and methods (Ashby et al., 2006; Caputi et al., 2006; Schael, 2009; Koprowski, 2015) (Section III.1.2). These approaches offer wider coverage and enable counterparts identification for marginal SCUBA-2 sources, which we classified as ‘SUPP’. Section III.1 outlines the feasibility and reliability of each identification method.
We show in Figure 7 the predicted flux densities as a function of redshift in various identification bands for a source that is just detected by SCUBA-2 (at mJy). Our identification methods are robust for sources out to at least –4. Note that the majority of SMGs have 850 fluxes around 4 mJy—i.e., brighter than the fiducial detection threshold used here—their multi-wavelength counterparts are expected to be even more easily detectable, enabling reliable identification out to higher redshifts. Moreover, the intrinsic diversity of SMG SEDs further enhances the potential of our approach to detect dusty galaxies during the epoch ending reionization (–6).
Several studies have selected sole counterparts based on the minimum -value criterion (Zavala et al., 2018), demonstrating that this approach typically recovers the majority of submillimeter flux (Simpson et al., 2015). However, we adopt a more conservative approach by retaining all potential counterparts.
By combining all identification methods, we matched 248 counterparts to 192 SCUBA-2 sources. The vast majority () of these identified galaxies have deblended 850 flux densities greater than 1 mJy. Notably, for the 212 newly detected SCUBA-2 sources (Section II.1), we successfully identified 136 SMG counterparts for 101 SCUBA-2 sources, representing more than 50% of the total confirmed counterparts. Figure 8 presents representative examples of multiwavelength counterpart identification. A more detailed breakdown of the identifications is provided in Section III.2.
We consider these methods to be generally robust and represent our best-effort approach. However, we emphasize that they are not perfect: some genuine SMGs are inevitably missed, while a fraction of spurious sources may also be included. We offer an in-depth discussion of the reliability of these identification methods and their potential limitations in Section III.3.
Table 2 presents the photometric data for the first ten SMG sources in our sample. The complete catalog, available in the online supplementary material, includes continuum photometry spanning from the optical/NIR to the radio bands; here we list only the measured flux densities in each band. Table 3 lists the first ten entries of our identified source catalog, including information on the SCUBA-2 sources, their galaxy counterparts, and derived physical properties. When a source is identified by multiple methods, we report only the most reliable counterpart according to the following hierarchy: . The two complete catalogs are available in the supplementary material.
=50mm {rotatetable*}
| id ID | id RA | id DEC | S | S | S | S aaThe sky-projected ACIS pixel size is 0.492 arcsec. | S aaThe far-infrared and (sub)millimeter flux densities are derived using the super-deblending technique, as described in Section II.4.2 and Section II.5. | S aaThe far-infrared and (sub)millimeter flux densities are derived using the super-deblending technique, as described in Section II.4.2 and Section II.5. | S | ||
| [deg] | [deg] | [Jy] | [Jy] | [Jy] | [Jy] | [mJy] | [mJy] | [mJy] | [mJy] | [Jy] | |
| SSA22.0000a | 334.385 | 0.295 | 2.250.42 | 8.950.06 | 10.475.63 | 10.871.35 | 6.792.50 | 37.462.96 | |||
| SSA22.0000b | 334.385 | 0.294 | 0.330.01 | 1.560.32 | 0.695.49 | 7.151.81 | 4.752.20 | ||||
| SSA22.0000c | 334.384 | 0.293 | 0.250.01 | 2.650.45 | 5.460.06 | 13.905.85 | 0.001.85 | 3.861.76 | |||
| SSA22.0001a | 334.328 | 0.301 | 0.100.01 | 4.130.55 | 7.650.06 | 0.220.02 | 21.276.25 | 7.170.92 | 3.441.23 | 93.313.13 | |
| SSA22.0001b | 334.328 | 0.303 | 4.000.05 | 23.304.35 | 4.470.83 | 3.081.23 | 10.683.34 | ||||
| SSA22.0002a | 334.404 | 0.306 | 2.670.01 | 3.610.50 | 10.070.06 | 29.971.78 | 7.3910.80 | 21.454.17 | |||
| SSA22.0002b | 334.404 | 0.306 | 3.070.49 | 10.070.06 | 29.971.78 | 2.7511.45 | 4.711.63 | 3.011.29 | |||
| SSA22.0002c | 334.403 | 0.305 | 0.230.02 | 2.620.48 | 3.830.05 | 13.295.50 | 2.301.13 | 3.111.56 | |||
| SSA22.0003 | 334.250 | 0.178 | 0.650.01 | 27.706.81 | 11.091.26 | 5.591.45 | 122.1210.37 | ||||
| SSA22.0004 | 334.391 | 0.231 | 13.710.95 | 21.440.06 | 39.771.78 | 15.544.92 | 7.241.15 | 6.441.21 | 39.842.36 |
Note. — The first ten entries of the SSA22 SMG photometric source catalog; the full catalog is available in the supplementary material.
=30mm {rotatetable*}
| id ID | S aaFor some sources, the deblended fluxes are poorly constrained due to source blending or low signal-to-noise. | id method bbThe 3 count sensitivity limit. | -value | z-flag ccRedshift types include spectroscopic redshifts, EAZY photometric redshifts, and CIGALE SED-fitting redshifts. When multiple redshift values are available, we adopt the most reliable one according to the hierarchy: (Section IV.2.2). | log | log SFR | log | log | MS ddThe offset of SMGs from the main sequence, defined as MS = log10(sSFR/sSFR), where the main sequence relation is adopted from Speagle et al. (2014). See also Paper IV. | ||
| [mJy] | [M⊙] | [M⊙ yr-1] | [L⊙] | [M⊙] | [mag] | ||||||
| SSA22.0000a | 10.87 | alma() | 0.02 | 3.09 | spec | 11.93 | 2.84 | 12.95 | 9.78 | 3.97 | -0.27 |
| SSA22.0000b | 7.15 | 0.04 | 0.59 | eazy | 8.97 | 1.76 | 11.56 | 9.95 | 5.78 | 1.70 | |
| SSA22.0000c | alma | 0.05 | 3.09 | spec | 10.85 | 3.04 | 12.83 | 8.94 | 2.79 | 0.79 | |
| SSA22.0001a | 7.17 | alma() | 0.01 | 3.07 | spec | 10.71 | 3.25 | 13.08 | 9.19 | 3.15 | 1.12 |
| SSA22.0001b | 4.47 | alma() | 0.04 | 3.10 | cigale | 11.91 | 2.55 | 12.89 | 9.00 | 5.09 | -0.53 |
| SSA22.0002a | alma() | 0.00 | 0.58 | spec | 7.99 | 0.82 | 10.59 | 9.35 | 2.32 | 1.39 | |
| SSA22.0002b | 4.71 | alma() | 0.01 | 3.09 | spec | 11.73 | 2.64 | 12.75 | 8.62 | 3.17 | -0.31 |
| SSA22.0002c | 2.30 | alma | 0.02 | 3.09 | spec | 10.71 | 1.94 | 12.89 | 7.77 | 1.84 | -0.21 |
| SSA22.0003 | 11.09 | 0.00 | 0.56 | eazy | 9.59 | 1.80 | 11.64 | 10.04 | 5.10 | 1.37 | |
| SSA22.0004 | 7.24 | 0.01 | 2.56 | spec | 10.98 | 2.79 | 12.81 | 9.63 | 2.00 | 0.52 |
Note. — The first ten entries of our identified source catalog, the full catalog is available in the supplementary material.
III.1 Identification Methods
III.1.1 Radio+ Methods
Radio and mid-infrared (MIR) emission serve as reliable tracers of star formation activity in galaxies, exhibiting strong correlations with submillimeter emission.
Due to the radio-FIR correlation spanning several orders of magnitude in luminosity (Condon, 1992), it remains well-preserved even at high redshifts. Additionally, with its relatively low spatial density (Koprowski, 2015), high-precision radio interferometric observations are widely used for counterpart studies of submillimeter sources (Dudzevičiūtė et al., 2020b). Although the origin of the radio-FIR correlation is debatable, it is widely believed to trace the activity of massive stars () in galaxies (Helou et al., 1985; Condon, 1992; Dudzevičiūtė et al., 2020b). Submillimeter radiation arises from the re-emission of ultraviolet photons absorbed by dust from star-forming activities in galaxies (Blain, 2002). Radio emissions in galaxies arise from synchrotron and free-free radiation. Due to their flat spectrum, free-free emissions are weaker than synchrotron emissions at wavelengths longer than 1 cm (frequencies 30 GHz). Therefore, radio emissions are generally attributed to the synchrotron acceleration of relativistic electrons produced by HII regions of massive stars and the supernova remnants left after their death (Condon, 1992).
Similar to the former, the MIR emissions also exhibit a strong correlation with FIR emissions (Appleton et al., 2004; Takeuchi et al., 2005), serving as a reliable indicator of the infrared luminosity of dust-obscured galaxies (Toba et al., 2017; Shim et al., 2022). A significant portion of the MIR emissions from SMGs originates from thermal emissions, generated by dust grains heated by star-forming regions, AGN, or nuclear starburst regions. Moreover, MIR emissions also stem from atomic and molecular lines, primarily induced by the bending and stretching of polycyclic aromatic hydrocarbons (PAHs), excited by UV photons associated with nearby star formation (Menéndez-Delmestre et al., 2009; Casey et al., 2014; Boquien et al., 2019). The dominant PAH feature lines are emitted at rest-frame wavelengths of 3.3, 6.2, 7.7, 8.6, 11.3, 12.7, and 17 (Draine and Li, 2007; Dudzevičiūtė et al., 2020b), falling within the observed 24 band at typical redshifts of SMGs. We should note that it is precisely due to these prominent PAH emission and absorption features (Huang et al., 2009) that the 24 flux in SMGs can exhibit significant variations.
We employ the corrected-Poissonian probability, -value, to calculate the probability of chance matches between the identified counterparts and submillimeter sources, as outlined by Downes et al. (1986):
where is the surface density of VLA 3 GHz (or 24 ) sources, and is the angular distance between identifications and the SCUBA-2 source positions.
Hodge et al. (2013) argued that, even in single-dish submillimeter maps where source confusion is significant, the positional offset between the submillimeter source and true SMG counterpart does not appear to vary systematically with SNR. An et al. (2018) adopted a statistical threshold of 0.065 for reliable radio identifications, though they noted that a small fraction of confirmed radio counterparts still lie beyond this threshold—corresponding to 5.5. Similarly, Chen et al. (2016) found that using a fixed search radius for counterpart identification could be a better strategy, and suggested that adopting 0.1 balances accuracy and completeness.
We therefore adopt a fixed search radius of 7 (half of the SCUBA-2 beam size) and consider all VLA and MIPS 24 counterparts with 0.1 as potential associations. Specifically, all VLA sources within a 5.4 radius of a SCUBA-2 centroid have 0.1. For example, in Figure 9, the sources SSA22.0000, SSA22.0070, and SSA22.0002 are affected by source blending. Our identification method did not associate the radio counterparts located near the edge of the SCUBA-2 beam with these sources; however, ALMA detections confirm that these radio sources are indeed physically associated SMGs. For MIPS 24 , due to its lower source surface density, all counterparts within the full SCUBA-2 beam (7) satisfy 0.065. Taking into account SCUBA-2 beam effect, source blending, uncertainties in source extraction, astrometric errors, and telescope pointing accuracy ( 1–2 for JCMT; Geach et al., 2017), we regard the 0.1 criterion as well justified for radio counterparts. Given that 24 band has a resolution of 6, we similarly consider this threshold appropriate for 24 identifications. For the IRAC 8 and -color identifications used in the next section, we adopt a stricter threshold of 0.05, corresponding to angular radii of approximately 2.4 and 2.8, respectively. We provide a more detailed justification for these choices in Section III.3.
For convenience, we will continue the tradition of labeling sources in the catalog as ‘robust’ for 0.05 and ‘tentative’ for 0.05. After visually inspecting all counterparts and candidate sources, we defined sources with coordinate separations less than 1.5 as the same counterpart. This threshold is appropriate as it avoids misclassifying sources that are too far apart as the same, or splitting the same source into different counterparts. Multiple identifications for the same SCUBA-2 source were sorted in ascending order of -values and assigned labels a, b, c, and so forth. Additionally, we found that despite the larger FWHM of 24 sources compared to high-precision VLA detections, their coordinates are still relatively accurate.
III.1.2 + Methods
We also employ IRAC 8 and red opt-NIR color as a method to search for potential counterparts. However, for reliability considerations, we only include those counterparts with 0.05 in the source catalog.
Similar to the MIR-FIR correlation mentioned earlier, exploration in the IRAC 8 band is often conducted to enhance the completeness of identifying submillimeter sources (Ashby et al., 2006; Koprowski, 2015; Koprowski et al., 2017; Zavala et al., 2018). SED analysis of SMGs suggested that their 8 emission primarily traces ongoing star formation (Ashby et al., 2006). The evolved stellar populations have been discovered in massive SMGs, the expected spectral signatures of evolved stellar populations should peak at rest-frame 2-3 (Koprowski, 2015), corresponding to observed 8 .
Some studies have used a NIR color to select high-redshift, extremely red galaxies, such as HIEROs (Wang et al., 2012), KIEROs (Wang et al., 2016), OIRTC (Chen et al., 2016). Similarly, due to the highly dust-obscured nature of SMGs often displaying red optical/NIR colors, many researchers have used color-selected galaxies to identify DSFGs or SMGs (Cowie et al., 2017; An et al., 2018; Zhang et al., 2022; McKay et al., 2025). Following Michałowski et al. (2012), we employed a red optical/NIR color criteria ( 2 & 23 mag) to identify potential SMGs counterparts. The detailed measurement description could be seen in the Section II.3.
SMGs typically exhibit redshifts in the range –3. In this redshift interval, the observed - and -bands correspond approximately to the rest-frame near-UV and -band, respectively, straddling the Balmer/4000Å break—a key spectral feature for galaxy characterization. The Balmer break arises at the Balmer limit (3645 Å) due to bound-free absorption by hydrogen atoms with electrons in the energy level. This feature is weakened in hot OB stars because most hydrogen is ionized, but it is strongest in cooler A-type stars (– M⊙), whose effective temperatures (10,000 K) maximize the population of electrons in the state. This feature is particularly prominent in star-forming galaxies (Dudzevičiūtė et al., 2020b; Wilkins et al., 2023). Moreover, it becomes especially strong in galaxies with stellar populations of age 0.3 Gyr, when a large number of UV-bright stars begin to evolve off the main sequence (Bruzual and Charlot, 2003; Vikaeus et al., 2024). The 4000 Å break primarily driven by metal absorption lines from cool, evolved stars (Mo et al., 2010; Dudzevičiūtė et al., 2020b). Together, the Balmer/4000Å break serves as a powerful diagnostic of recent star formation history and is sensitive to dust attenuation, metallicity, and the Lyman continuum (LyC) escape fraction (Wilkins et al., 2023).
These selected galaxies are referred to as Extremely Red Galaxies (ERGs), which may stem from evolved stellar components or galaxies that have reddened significantly due to dust obscuration in star-forming galaxies or starbursts (or AGN) (Caputi et al., 2006; Schael, 2009; Mo et al., 2010). ERGs are believed to align on an evolutionary sequence with high-redshift SMGs and local massive ellipticals (Smail et al., 2002; Schael, 2009), while the most extreme ERGs represent a population of highly dust-obscured, ultraluminous galaxies at high redshift (Smail et al., 1999). In the redshift range of 1.5-3, 40-60% of high-mass ERGs are ULIRGs (Smail et al., 2002; Caputi et al., 2006). SMGs, characterized by severe dust obscuration and intense star formation, are expected to exhibit distinct optical/NIR features. Therefore, such selections seem to be a good criterion.
III.2 Identification Result
Our VLA 3 GHz and MIPS 24 catalog includes 1989 radio sources with a significance greater than 4 (the image depth after primary beam correction is 1.5-5 Jy), and 813 sources of 3 at 24 , respectively.
Armed with the robust methods, we identified a total of 173 SMG candidates for 142 SCUBA-2 sources, with 36 candidates detected by both 3 GHz and 24 . These sources are predominantly located in the main SCUBA-2 map area of SSA22 (Figure 1), including the protocluster core (Zeng et al., 2024).
Specifically, of the 203 SCUBA-2 sources whose positions lie within the area covered by VLA 3 GHz observations, 119 have VLA counterparts, yielding an identification rate of 59%. In total, the VLA data reveal 146 SMG counterparts across the full sample. On the other hand, 218 SCUBA-2 sources are located within the MIPS 24 map, with 63 SCUBA-2 sources having 24 counterparts, resulting in an identification rate of 29%. Among the 63 SCUBA-2 sources, there are 63 SMG counterparts at 24 sources, 27 of these counterparts are additional findings not captured by the VLA observations. The overall identification rate for the + method is 56%. Our SCUBA-2 and VLA 3 GHz maps are similar in depth to those of Hyun et al. (2023), and our radio-identification rate is approximately the same.
The ADF22 catalog confirmed 20 SMG counterparts associated with 15 SCUBA-2 sources. Among these SMGs, 5 were not identified by either the or methods, though 2 of those 5 were also detected as IRAC 8 sources. Additionally, one ALMA-detected SMG serves as the unique counterpart to the SCUBA-2 source SSA22.0098.
In total, utilizing the radio+MIR method, we obtained 178 reliable SMG counterparts for 143 SCUBA2 sources, identified by ALMA, VLA 3 GHz, and MIPS 24 sources.
96 candidates were identified by 8 sources, while 93 by the optical/NIR-red selected galaxies. These two approaches respectively added 30 and 40 new SMG counterparts, constituting 28% of the total identifications. Among these, 18 and 29 SCUBA-2 sources were uniquely identified with counterparts by the and methods, respectively.
III.3 Identification Accuracy and Caveats
In this section, we provide a preliminary assessment of the reliability of our identification methods and offer important caveats for users of our catalog.
To evaluate the performance of our identification techniques, we cross-match our identified sources with two ALMA surveys in the SSA22 field: the ADF22 survey (including the ADF22 A/B fields; Umehata et al., 2017, 2018) and the deeper ADF22+ survey (Huang et al., 2025), as well as the ALMA 2 mm continuum survey (Cooper et al., 2022).
Figure 9 shows the footprint of the ADF22 survey. Focusing on SCUBA-2 sources detected at 4, the identification accuracies of our four methods—, , , —are 15/20 ( 75%), 5/6 ( 83%), 9/9 ( 100%), and 2/3 ( 67%), respectively. When including marginal detections in the range 3.5–4, the accuracies decrease modestly to 17/27, 5/6, 10/12, and 2/4. The high fidelity for 4 sources reflects the fact that the ADF22 fields lie within the region of most complete multiwavelength coverage, underscoring the critical importance of deep, multi-band data for reliable source identification. Survey depth also plays a role: the deeper ADF22+ observations confirmed three low-significance ADF22 A detections, one of which (ADF22+ source A28) corresponds to a component of the blended SCUBA-2 source SSA22.0070. This source was robustly identified by both our and methods.
Comparison with the ALMA 2 mm survey yields (Figure 10) lower—but still reasonable—identification accuracies: 13/19 ( 68%), 2/5 ( 43%), 9/12 ( 75%), and 8/13 ( 62%). We emphasize again that these accuracy estimates depend sensitively on both multiwavelength coverage and ALMA survey depth.
We further compare sources identified via and methods with those detected in radio and 24 . The matching fractions are consistent between the two methods (Figure 12), reaching 75–85% in the central region of the field (where SCUBA-2 map RMS 1 mJy). Both matching rates decline with increasing SCUBA-2 map noise, yielding overall matching fractions of 71% () and 67% ()—a direct consequence of the deepest and most complete multiwavelength coverage in the field center. Because IRAC 8 sources and color-selected objects have relatively high surface densities, positional association probabilities drop below 0.05 only at angular separations of 2.4–2.8. The median offsets between our and identifications and the SCUBA-2 centroids are 1.5 and 1.8, respectively—corresponding to physical scales of 10–15 kpc at 2. This is comparable to the pointing uncertainty of JCMT (Geach et al., 2017). Therefore, if SMGs reside near the centroid of the SCUBA-2 source, these two identification methods are unlikely to introduce significant misidentifications at such angular separations, aside from chance alignments along the line of sight.
The the 80% correspondence between sources identified via the / methods and those detected in radio or is remarkably high—indeed, a pleasantly surprising result. This strong overlap suggests an underlying physical connection: even when some of the IRAC 8 sources or color-selected objects identifications do not correspond to genuine SMGs, they are still likely to be galaxies undergoing active star formation. Moreover, as shown in the earlier matching analysis, the identification accuracy of the 8 method reaches 80%, demonstrating its considerable potential—comparable to the more established and techniques.
For the method, we perform additional tests to quantify the contamination from random alignments. First, within the ADF22 survey area, we mask all regions within 7 of any ALMA-detected SMG or SCUBA-2 source. We then place random positions and check whether a source lies within 2.78 ( 0.05). This yields a lower limit on the contamination rate due to chance superpositions. In this small masked region, we generate 200 random positions per trial and repeat the experiment 50 times, obtaining a false-match probability of 4.02 1.32%. Next, we expand the test to a larger rectangular area (R.A. = –, Dec. = +000700–+002800), which covers field center with the most complete multiwavelength coverage. Here, we draw 2000 random positions per trial over 10 repetitions, resulting in a false-match probability of 4.84 0.40%. Overall, thanks to our strict matching criterion ( 0.05), the expected contamination from random alignments is low—below 5%. However, it should be emphasized that this estimate represents a lower limit. In reality, the true contamination could be higher, because SCUBA-2 sources are preferentially located in overdense environments. Such clustering enhances the likelihood of finding a red, massive galaxy near a submillimeter position even without a physical association.
Deep, well-characterized ALMA follow-up surveys—such as ALESS (cunha2015) and AS2UDS (Stach et al., 2019; Dudzevičiūtė et al., 2020b)—contain only a small minority of sources at low redshift ( 1), with the vast majority lying on or above the star-forming main sequence. In contrast, our sample includes a non-negligible fraction of low-redshift or quiescent systems, regardless of the identification method used (see Figure 11). We acknowledge, however, that some of these apparently quiescent galaxies may be artifacts of SED fitting: When far-infrared constraints are weak or absent, it leads to increased scatter in derived physical properties and may result in spurious classifications as “quenched” systems. We briefly discuss this effect in Section V.2. We also emphasize that our multiwavelength identification approach inevitably leads to the loss of a fraction of high-redshift SMGs. To illustrate this, we apply simple selection criteria to the AS2UDS catalog (Dudzevičiūtė et al., 2020b), isolating subsamples with radio detections, MIPS 24 counterparts, IRAC 8 counterparts, or those satisfying the color cut 2 and 23 mag. The median redshifts of these subsamples are 2.45, 2.31, 2.58, and 2.29, respectively—systematically lower than that of the full AS2UDS sample ( 2.61). This trend underscores a key limitation: multiwavelength identification methods inherently risk missing the high-redshift SMGs. This bias is particularly pronounced when relying on 24 or optical/NIR data. In the former case, the incompleteness is likely driven by sensitivity limits at high redshift, while in the latter, severe dust attenuation renders many high-redshift SMGs undetectable in optical/NIR bands.
We issue the following cautions: (1) Sources in regions with incomplete multiwavelength coverage—approximately where the SCUBA-2 map RMS 1–1.2 mJy—should be treated with care. (2) Identifications based solely on the criterion require particular scrutiny due to higher contamination risk. (3) Sources with redshifts 1 or classified as quiescent are likely contaminants and should be interpreted cautiously. Secondly, it is important to emphasize that the current identification methods likely miss a fraction of genuine SMG counterparts, particularly at high redshifts. Lastly, while our identification methods recover the majority of the SCUBA-2 flux, this raises a secondary concern: misidentified counterparts could lead to incorrect deblending and flux allocation. While given the relatively low multiplicity rate and the small number of -only identifications, we consider this risk to be modest.
In summary, cross-matching with deep ALMA data indicates that the majority of our identified sources belong to the true SMG population. Consequently, our sample is suitable for robust statistical analyses—a conclusion we will reiterate in subsequent sections.
III.4 Multiplicity
ALMA interferometric observations have revealed that single-dish submillimeter sources are often blends of multiple SMGs (Karim et al., 2013). In a follow-up ALMA study of 716 UDS SCUBA-2 sources, Stach et al. (2018) found an overall multiplicity fraction of approximately 11%. This fraction increases with submillimeter flux density: it rises to 26% for sources with S 5 mJy and further to 44% for those brighter than 9 mJy. Among these blended systems, about 30% consist of physically associated components (Stach et al., 2018). The measured multiplicity also depends on observational depth, as fainter SMGs are commonly found to be blended even in ALMA “blank” maps (simpson2014). When such faint blends are included, the multiplicity fraction may increase by an additional 10–20% (Hodge et al., 2013; Simpson et al., 2015).
In SSA22 field, the overall multiplicity rate is %. In the “deep region” (Figure 1), where multiwavelength data are most complete, fraction rises to 4%, consistent with typical multiplicity rates reported in other multiwavelength identification studies ( 20–40%; Chen et al., 2016; An et al., 2019). Specifically, for SCUBA-2 sources with flux densities 8 mJy, the multiplicity rate is 24 3%. In contrast, for sources 8 mJy, the multiplicity rate increases significantly, reaching 47 13%. Within individual high-flux bins, the multiplicity fractions range from 40% to 60%, though these estimates carry large uncertainties due to small number statistics.
IV SED-fitting
We present the photometric redshifts derived from EAZY (Section IV.1), a description of our CIGALE configuration (Section IV.2.1), and the resulting SED fitting outputs (Section IV.2.2).
IV.1 EAZY Photometric Redshift
We employed the Python version of EAZY (Brammer et al., 2008) to estimate optical/near-infrared (-band to -band) photometric redshifts (photo-). Additionally, we adopted an updated template set named corr_sfhz_13 777https://github.com/gbrammer/eazy-photoz/tree/master/templates/sfhz, which incorporates redshift-dependent star-formation histories. The transmission curves for each bandpass were obtained from the SVO filter profile service (Rodrigo et al., 2012, 2024). The program combined and scaled several spectral templates to fit the observed fluxes of identified sources, iterating three times to correct zeropoint offsets (Straatman et al., 2016; Weaver et al., 2022) and optimize the photometric redshift. We adopted the output parameter ‘z_ml’, representing the maximum likelihood redshift, as the final redshift determination for the counterparts. The 16th and 84th percentile probability distribution of photo- were used to calculate the 1 error.
To assess the reliability of our photometric redshifts, we performed a cross-match between the Subaru/Scam -band catalog and with spectroscopically confirmed galaxies (the “representative spectroscopic redshift catalog” compiled by Mawatari et al. 2023) within 1, yielding a sample of 1734 optically selected galaxies for photometric redshift validation.
Our analysis demonstrates that photometric redshifts achieve reliable accuracy when computed using at least 4 photometric bands, with 1645 out of the 1734 -selected galaxies meeting this criterion for valid EAZY redshift estimates. These validated sources are represented by gray points in Figure 14, showing excellent agreement between photometric and spectroscopic redshifts. The ensemble exhibits a median redshift residual () of -0.013 ± 0.001 (uncertainty estimated via bootstrap resampling) with a standard deviation of 0.39. The method demonstrates strong consistency, with 80.5% and 76.5% of sources falling within redshift residual range of 0.2 and 0.15 respectively, confirming the robustness of EAZY photometric redshifts for our sample.
Among the 248 identified SMGs, 27 sources have reliable ‘’ and ‘’ (red circles in Figure 14). Their median redshift residual is -0.014 ± 0.022, with a standard deviation of 0.20. Specifically, 77.8% (74.1%) of sources show redshift residuals smaller than 0.2 (0.15). The comparison also shows good agreement between photometric and spectroscopic redshifts.
In the end, 155 identifications (at least 4 bands available) possess ‘’ values ( 62.5%), and the median redshift is 1.8 when considering both ‘’ and ‘’ collectively. Our median optical/NIR photo- was noted to be lower than the redshift distribution of SMGs, with over a third of the counterparts lacking sufficient detections in the optical/NIR bands. Among these sources, 41 have not been detected at all. This observation suggests that the characteristics of this population are situated in the high-redshift universe and/or are highly dust-obscured.
IV.2 CIGALE SED Fitting
We perform SED fitting using Investigating GALaxy Emission (CIGALE version 2022.1; Noll et al., 2009; Boquien et al., 2019). By leveraging multi-wavelength photometry—from the optical/NIR through to radio—we derive key physical properties of our SMG sample. For galaxies lacking reliable or (Section IV.1), we additionally obtain their far-infrared (FIR) photometric redshifts. This subsection describes the fitting configuration, and the resulting physical parameters.
CIGALE adopts a modular design, facilitating the invocation and combination of different physical component models within galaxies, including stellar populations, star formation histories, dust emission and attenuation, radio emission, AGNs, nebular emission, and more. The SED fitting process employs an energy balance approach, where the stellar and AGN emission absorbed by dust in the UV/optical regime is re-radiated in the infrared through self-consistent dust modeling (Casey et al., 2014). CIGALE generates comprehensive SEDs for galaxies by exploring all possible combinations of model parameters through an extensive grid (typically containing millions to billions of parameter sets). The optimal SED fit is determined by calculating and minimizing the reduced statistic for each template.
The multiple galaxy factors, such as stellar age, metallicity, and attenuation often exhibit degeneracies (Boquien et al., 2019), which can lead to galaxies with diverse properties potentially exhibiting similar SEDs. Therefore, CIGALE employs a Bayesian-like approach that weights all templates based on the goodness of fit to estimate physical properties and uncertainties, as well as partially mitigates the degeneracies effects between physical parameters (Boquien et al., 2019).
| Modules | Description | Parameters | Values |
| SFH: sfh2exp | age of the main stellar population | age | 500, 1000, 3000, 5000, 10000 |
| e-folding time of the main stellar population | tau_main | 1000, 3000, 5000, 10000 | |
| age of the burst population | burst_age | 10, 50, 100, 300 | |
| e-folding time of the burst population | tau_burst | 30, 100, 300, 9000 | |
| mass fraction of the burst population | f_burst | 0, 0.01, 0.1, 0.3, 0.5 | |
| SSP: bc03 | initial mass function | imf | 1 (Chabrier (2003)) |
| metallicity | 0.02 | ||
| Nebular emission: nebular | default | default | |
| Dust attenuation: | V-band attenuation in ISM | Av_ISM | 0.3, 0.6, 1.0, 1.6, 2.3, 3.0, 3.8, 5.0, 8 |
| dustatt_modified_CF00 | Av_ISM / (Av_BC+Av_ISM) | mu | 0.44 |
| attenuation power-law slope in the ISM | slope_ISM | -0.7 | |
| attenuation power-law slope in the BC | slope_BC | -1.3 | |
| Dust emission: dl2014 | mass fraction of PAH | qpah | 0.47, 1.12, 2.50, 3.90, 6.63 |
| minimum radiation field | umin | 0.1, 0.5, 1.0, 5.0, 10.0, 25.0, 50.0 | |
| radiation field power-law slope | alpha | 2.0, 2.5, 3.0 | |
| AGN: fritz2006 | angle between equator and line of sight | psy | 0.001, 50.1, 89.990 |
| AGN fraction | fracAGN | 0.0, 0.001, 0.01, 0.1, 0.3, 0.5, 0.7, 0.9 | |
| optical depth at 9.7 | tau | 1.0, 6.0 | |
| opening angle of the dust torus | opening_angle | 100.0 | |
| polar dust temperature | temperature | 100.0 | |
| emissivity index of the polar dust | emissivity | 1.6 | |
| Radio: radio | FIR/radio correlation coefficient | qir_sf | 2.15, 2.45, 2.78 |
| power-law slope of emission related to stellar | alpha_sf | 0.8 | |
| power-law slope of AGN radio emission | alpha_agn | 0.7 | |
| redshifting | redshift grid | redshift | 0.0001, |
| 0.1 to 1.4 with a step of 0.2, | |||
| 1.5 to 3.9 with a step of 0.1, | |||
| 4 to 6 with a step of 0.2 |
IV.2.1 Configuration
The CIGALE configuration models and parameter settings adopted in our work are summarized in Table 4.
We adopt the sfh2exp model with a double exponential star formation history (SFH). By setting the mass fraction of the second burst, we determine whether the model experiences one or two star formation events, with the latter including one long-term star formation and one recent burst. Previous studies suggest that SMGs may undergo multiple intense star formation episodes, and an exponentially declining SFH also consistent with true values (Michałowski et al., 2014). Both simulations and main-sequence models indicate that double SFHs provide a more likely description, regardless of whether galaxies evolve through mergers or in isolation (Speagle et al., 2014).
We employ the initial mass function (IMF) model from Chabrier (2003) and adopt the simple stellar population (SSP) templates from Bruzual and Charlot (2003). The SSP models incorporate all phases of stellar evolution from zero-age main sequence to thermally pulsating asymptotic giant branch (TP-AGB) stages for low- and intermediate-mass stars, and core-carbon ignition for massive stars. The stellar models effectively constrain galaxy properties including star formation history, metallicity, and dust mass.
The continuum emission from thermally ionized gas in star-forming regions (e.g., HII regions) spans optical to radio wavelengths and may represent a non-negligible component for high-redshift galaxies (Boquien et al., 2019). We implement the nebular emission template from (Inoue, 2011).
Dust plays a fundamental role in the energy balance approach. For dust attenuation, we employ the physically motivated dusttt_modified_CF00 model, which accounts for dust grain mixing and geometric effects while distinguishing between different stellar environments (birth clouds and interstellar medium (ISM)). Tests on infrared-selected galaxy samples demonstrate that this model more accurately reproduces dust attenuation properties (Lo Faro et al., 2017; Barrufet et al., 2020).
The dl2014 templates, an updated version of Draine and Li (2007), are adopted to model dust emission, which incorporates a mixture of amorphous silicate and graphite grains with variable PAH abundances. The dust emission consists of two components: diffuse dust heated by the general interstellar radiation field and localized emission from star-forming regions. Analysis of 850 sources by Seo et al. (2018) indicates that dl2014, with its more comprehensive physical parameters, provides superior fitting results. The 850 selection preferentially picks out colder SMGs, whose ISM emission peaks at longer wavelengths. Accordingly, we include lower radiation field intensities () in our parameter space.
AGNs may play a crucial role in driving galaxy evolution (Toft et al., 2014), although not all SMGs necessarily undergo an AGN phase (Johnson et al., 2013). Observations indicate that 20-40% of SMGs show signs of AGN activity (Alexander et al., 2005a; Laird et al., 2010; An et al., 2019; Stach et al., 2019). In systems with strong AGN contributions, the radio emission from the AGN can compete with or even dominate over that from star formation, necessitating the inclusion of an AGN component in our analysis. We adopt the AGN model from (Fritz et al., 2006), which self-consistently treats the central active nucleus, dust scattering in the torus, and thermal re-emission while accounting for viewing angle effects. Our fitting experience reveals that the optical emission in many sources cannot be fully explained by dust-attenuated stellar radiation alone, often showing a smooth transition to a “plateau” at bluer wavelengths (e.g., Figure 16, sources SSA22.0026a, SSA22.0053, and SSA22.0065). The inclusion of even minimal AGN contributions (fractions of 0.001-0.01) significantly improves the SED fits by better reproducing this blue excess emission.
The radio emission in most galaxies is predominantly powered by star formation activity, consisting of the non-thermal synchrotron radiation from relativistic electrons accelerated in supernova remnants, and thermal bremsstrahlung emission (free-free emission) from HII regions surrounding massive stars. While bremsstrahlung exhibits a flat spectrum, its contribution becomes negligible compared to synchrotron emission at frequencies below 30 GHz (Helou et al., 1985; Condon, 1992). The well-established radio-FIR correlation arises because both wavebands trace star formation activity (Section III.1.1). However, we allow for variations in the correlation coefficient to account for potential redshift evolution of the radio-FIR relation (Speagle et al., 2014), and possible radio excess caused by AGN contamination.
For sources lacking spectroscopic redshifts or those too optically faint for reliable EAZY photometric redshift estimates, we simultaneously derive their redshifts during the CIGALE SED fitting process. To maintain computational efficiency and optimize parameter space exploration, we adopt redshift bins with varying resolution: coarse sampling ( = 0.2) for low/high-redshift range ( 0-1.5 and 4-6), fine sampling ( = 0.1) for the critical redshift range ( 1.5-4), where the majority of SMGs are expected to reside (Dudzevičiūtė et al., 2020b).
IV.2.2 Fitting Result
For sources lacking ‘’ and ‘’ estimates, we derive their redshifts () simultaneously during the CIGALE SED fitting process using the photometric redshift mode.
Comparison with and reveals residuals of 0.091 ± 0.053 and 0.052 ± 0.022, respectively (Figure 15). Given the typically higher accuracy of photometric redshifts derived from optical/NIR bands compared to longer wavelengths, we adopt the , if available, followed by the , and then the in analysis.
We conducted visual inspections of all SED fitting results, excluding sources with poor fits. Specifically, we removed sources with weak or no constraints in the FIR/submillimeter bands, as well as those lacking constraints in other critical wavelength ranges. This quality control process eliminated 27 sources ( 10% of the sample) from our final analysis.
Additionally, we identified six sources (SSA22.0113, SSA22.0193, SSA22.0196a, SSA22.0209, SSA22.0329, and SSA22.0374) that consistently produced poor fits when using their fixed photometric redshifts. Our analysis suggests potential inaccuracies in the EAZY photometric redshifts for these sources. Allowing their redshifts to vary as free parameters yielded significantly improved fits, and we consequently adopted these revised solutions (marked in blue in the relevant figures).
Our final catalog contains 221 SMGs (corresponding to 186 SCUBA-2 sources) after excluding 27 unreliable fits and incorporating the 6 revised solutions. Notably, this includes 115 newly detected and analyzed SMGs associated with 95 SCUBA-2 sources, representing important additions to the known population of submillimeter galaxies.
V RESULTS
In this section, we present and discuss the physical properties of our identified sample, including redshift (Section V.1), SFR/stellar mass (Section V.2), infrared luminosity (Section V.3), dust mass (Section V.4), and dust attenuation (Section V.5).
liao2024; uematsu2024 compared the physical properties derived from the CIGALE and MAGPHYS. They found that differences between dust modeling assumptions in these codes lead to modest systematic offsets in estimated dust masses and temperatures, as well as in stellar masses. Our CIGALE configuration closely follows that of liao2024; therefore, we expect good overall agreement between our derived physical parameters and those from MAGPHYS—except for dust mass, which we explicitly discuss in Section V.4. In the following sections, we will directly compare our estimates with published values.
V.1 Redshift Distribution
Our final sample comprises 221 SMGs (counterparts to 186 SCUBA-2 sources), with a median redshift of 2.00 0.08 (bootstrap error) and 16th-84th percentiles of 0.97-3.08. Our results are consistent with previous multiwavelength identifications of SMGs, which typically report median redshifts of 2-2.3 (Zavala et al., 2018; An et al., 2019; Barrufet et al., 2020; Hyun et al., 2023). However, larger samples of ALMA 870 -selected SMGs show systematically higher median redshifts of 2.5-2.7 (cunha2015; Cowie et al., 2017; Dudzevičiūtė et al., 2020b). As noted by Hyun et al. (2023), the redshift distribution of SMGs identified through radio and near-infrared detection appears truncated at 4, resulting in a lower median redshift of 2, a finding that aligns well with our results. This discrepancy highlights the limitations of multiwavelength identification methods, where finite survey depths and uneven wavelength coverage can lead to incomplete identification of high-redshift SMGs.
Figure 17 panels (a) and (d) show the redshift histogram and spatial distribution of the full SMG sample, respectively. The horizontal bars in panel (a) represent the comoving cosmological volume probed within each redshift bin. The data reveal a pronounced evolution in SMG number density: at z 4, the density increases sharply by approximately a factor of 6, while at 1-3 it rises more gradually, maintaining values between 1.78-3.16 10-5 cMpc-3. This two-phase evolutionary pattern suggests distinct regimes of SMG formation and growth across cosmic time (see Paper IV of this series). Notably, this includes 115 newly identified SMGs (corresponding to 95 SCUBA-2 sources with fluxes 2.8-3.5 mJy) that we report for the first time, which exhibit a slightly lower median redshift of 1.9.
Our analysis reveals a clear dichotomy in the redshift distribution based on optical/NIR detectability: approximately two-thirds of the SMGs in our catalog qualify as optically bright sources (with 4 available optical/NIR photometric bands; cunha2015), exhibiting a median redshift of 1.75, significantly lower than the typical SMG redshift distribution. The remaining one-third, classified as optically faint sources ( 4 available bands), show a higher median redshift of 2.2. Moreover, the 27 sources completely undetected in optical/NIR bands (excluding SSA22.0347 which lies outside the optical/NIR images coverage) demonstrate an even more extreme redshift distribution, with a median of 3.0. This value not only exceeds that of the optically faint population but also lies above the characteristic redshift peak for SMGs. These results confirm that SMGs represent a population of heavily dust-obscured systems preferentially located in the high-redshift universe, with the most severely obscured members tending to reside at even greater cosmological distances.
When multiple redshift estimates are available for a source, we adopt them in order of reliability: (Section IV.2.2). Our sample contains 30 sources with spectroscopic redshifts (), 115 with EAZY photometric redshifts (), and 76 relying solely on CIGALE long-wavelength photometric redshifts (), exhibiting median redshifts of 2.42, 1.65, and 2.15, respectively.
The “MAIN” sample of 155 sources identified through radio and “” methods shows a median redshift of 2.07, with those having SCUBA-2 SNR 4 displaying a slightly higher median of 2.2. These results are consistent with the S2TDF survey under comparable SCUBA-2 depths and identification methods (Hyun et al., 2023). The supplementary (“SUPP”) sample of 66 sources identified solely via “” and “” methods exhibits a lower median redshift of 1.83 (Figure 17 image (b)). However, when considering all “/” selected galaxies (including those also detected in “radio/”), the median redshift rises to 2.00, matching the overall sample distribution.
When examining the results from each method individually (Figure 11): “” detection identifies 124 SMGs with median 2.10; “” selects 62 SMGs at median 1.92; “” recovers 94 SMGs with median 2.12; and the “” method identifies 91 SMGs with median 1.80.
The lack of significant differences in properties between the ‘MAIN’ and ‘SUPP’ samples (Table 5), as well as across identification methods, validating the “” and “” methods as effective tools for identifying high-redshift SMGs. The higher median redshift of -selected sources particularly highlighting the importance of observational depth for robust identification.
Cowie et al. (2017, 2018) reported that SMGs with brighter submillimeter fluxes tend to lie at higher redshifts. Analyses of the ALESS and AS2UDS samples by simpson2014; simpson2017 found a similar trend but argued that it is not statistically significant and may be an artifact caused by incomplete redshift information at fainter flux levels. Stach et al. (2019) demonstrated this phenomenon using a sample of 708 AS2UDS SMGs, attributing it to the “downsizing” effect (Cowie et al., 1996). Additionally, due to the negative -correction, higher flux densities tend to be found at higher redshifts, indicating significant infrared luminosity evolution. This trend is driven by the presence of higher gas fractions and stronger SFR in distant galaxies, resulting in greater luminosities (simpson2020; birkin2021). Dividing our SMGs into five flux bins, we find:
a slope consistent with AS2COSMOS, AS2UDS, and CO surveys (0.06–0.09 mJy-1; Stach et al., 2019; simpson2020; birkin2021) (Figure 18). In contrast, the median 850 flux shows no clear evolution with redshift, remaining near 4 mJy. This indicates that while the brightest SMGs favor higher redshift, the general SMG population shows no strong link between submillimeter flux and redshift.
V.1.1 Overdensity
C. C. Steidel, K. L. Adelberger, M. Dickinson, M. Giavalisco, M. Pettini, and M. Kellogg (1998); 113 identified a large-scale structure at in the central region of the SSA22 deep field using low-mass LAEs. Our identified SMGs also exhibit significant overdensity within this redshift slice (Figure 17 (c)).
When considering the broader redshift interval – the overdense region contains 17 SMGs, corresponding to a 8.4 (7.8) overdensity for bin widths of 0.05 (0.1). The ADF22 survey (Umehata et al., 2017, 2018) contributes five additional counterparts, two of which lie within the overdense region (green shaded area), confirming that our identification methods already robustly trace this large-scale structure at 7 significance.
Figure 17 plot (e) shows the spatial distribution of SMGs within the structure; nearly all reside in the dense core. Within the central pMpc2 region at —approximately corresponding to the “SSA22a” region (white dashed rectangle)—we find 13 SMGs. This yields a surface density of 1.3 pMpc-2, consistent with the AzTEC SMG density reported by Umehata et al. (2014). Assuming these galaxies reside within the cosmic web node traced by LAEs (with a characteristic depth of 4.9 pMpc; Matsuda et al. 2005), their comoving volume density is 0.27 pMpc-3 (13/47.5).
The 850 fluxes of these SMGs are fairly uniformly distributed between 1 and 10 mJy, peaking at 4–6 mJy. Their total stellar mass sums to M⊙, and their combined SFR reaches M. This corresponds to volume-averaged contributions of M in stellar mass density and 193 M in SFR density. The ultra-deep survey of the core region (ADF22+, 5; Huang et al., 2025) yields an even higher stellar mass and SFR volume density, reaching 1.2 M and 296 M⊙ yr-1 pMpc-3, respectively.
We further estimate the overdensity using angular surface densities. From the SCUBA-2 source perspective, these 13 SMGs originate from 9 distinct SCUBA-2 detections, yielding an angular density of 700 deg-2—comparable to the S2CLS field average of 570 deg-2 (Geach et al., 2017). However, as noted by Wang et al. (2025), our 9 sources are confined to an extremely narrow redshift slice (, i.e., pMpc), whereas S2CLS is sensitive to a much broader redshift range (e.g., –, or 1837 pMpc; Dudzevičiūtė et al. 2020b). Accounting for this, the number density in the overdense region is enhanced by a factor of 12 relative to the blank field.
Using deblended SMG catalogs for a fairer comparison, we compute the angular surface densities of SMGs in the same redshift window () from AS2UDS (Dudzevičiūtė et al., 2020b) and JWST-TDF (Hyun et al., 2023), finding 66 deg-2 and 34 deg-2, respectively. The overdensity represents a factor of 15–29 enhancement over the field. Thus, overall, the SMG overdensity exceeds blank field average by more than an order of magnitude.
As shown in the figure 17 (e), Lyman-alpha Blobs (LABs)—objects exhibiting extreme galactic activity (Geach et al., 2005, 2014)—are co-spatial with our SMG sample within the large-scale structure of SSA22. Moreover, studies have shown that SMGs are frequently associated with other populations of galaxies within large-scale structures (Tamura et al., 2009; Umehata et al., 2014; Wang et al., 2025) and co-evolve with these cosmic environments (Umehata et al., 2019). These results highlight the role of large-scale structures in driving galaxy evolution, while conversely underscoring the potential of massive dusty galaxies as tracers of cosmic structure formation—particularly for mapping mass distributions in high-redshift overdense regions.
=60mm {rotatetable*}
| Porperties | Total sample | New detections aaCounterparts of submillimeter sources detected in our deeper observations. | ‘MAIN’ bbIdentification method. Most sources with ALMA counterparts are also recovered by our identification method. If additional identifications exist, the next most reliable method—ranked as —is indicated in parentheses. Owing to the superior astrometric precision of ALMA, we adopt ALMA-derived coordinates for source positions whenever available. | ‘SUPP’ | Conservative Sample ccTo minimize contamination, we adopt a conservative sample selection, retaining only sources with redshift 1 that have or counterparts. | 2.0 | 2.0 | Opt/NIR faint ddThe optically faint subsample consists of identifications with detections in fewer than four opt/NIR filters, whereas the optically bright subsample includes sources detected in four or more opt/NIR filters. | Opt/NIR bright |
| 221 ids | 115 ids | 155 ids | 66 ids | 133 ids | 110 ids | 111 ids | 73 ids | 148 ids | |
| Redshift | 2.00 | 1.90 | 2.07 | 1.83 | 2.24 | 1.25 | 2.70 | 2.20 | 1.75 |
| log Mstar [] | 11.19 | 11.20 | 11.23 | 11.03 | 11.28 | 10.93 | 11.32 | 11.29 | 11.14 |
| log SFR [ yr-1] | 2.22 | 2.07 | 2.25 | 2.06 | 2.38 | 1.58 | 2.55 | 2.25 | 2.12 |
| log LIR []) | 12.35 | 12.25 | 12.40 | 12.32 | 12.51 | 11.91 | 12.69 | 12.42 | 12.31 |
| log Mdust [] | 9.29 | 9.22 | 9.22 | 9.48 | 9.22 | 9.35 | 9.19 | 9.35 | 9.27 |
| log Mstar/ [] | 0.21 | 0.20 | 0.21 | 0.21 | 0.20 | 0.29 | 0.12 | 0.34 | 0.17 |
| [mag] | 3.09 | 2.82 | 3.08 | 3.16 | 3.09 | 3.10 | 3.06 | 3.87 | 2.82 |
| log Agem [yr] eeThe mass-weighted ages of the galaxies, computed as (Noll et al., 2009), are derived from CIGALE. | 8.75 | 9.04 | 8.75 | 9.05 | 8.70 | 9.17 | 8.62 | 8.91 | 8.75 |
| log sSFR [yr-1] | -9.05 | -9.18 | -9.08 | -9.03 | -9.00 | -9.37 | -8.83 | -9.15 | -9.00 |
| MS ffThe offset of SMGs from the main sequence, defined as MS = (sSFR/sSFR), where the main sequence relation is adopted from Speagle et al. (2014). See also Paper IV. | -0.16 | -0.21 | -0.19 | -0.03 | -0.20 | -0.20 | -0.15 | -0.27 | -0.08 |
Note. — Sample grouping is based on various classification schemes. Average properties represent median values of individual parameters, with uncertainties representing the 16th-84th percentile range.
V.2 SFRs and Stellar Mass
The distributions of SFRs and stellar masses () for the 221 SMGs identified in the SSA22 deep field are presented in Figure 19, with the histogram shown to the right.
The overall sample exhibits a median SFR of M⊙ yr-1, with the 16th–84th percentile range spanning 18–703 M⊙ yr-1. Over 90% of the sources have SFRs exceeding 10 M⊙ yr-1, and the highest SFR reaches M⊙ yr-1. The median stellar mass is M⊙, with 68% of the sample falling within the range 1011 M⊙. Approximately 92% of the sources have stellar masses M⊙, and the most massive galaxy reaches M⊙.
The stellar mass distribution of SMGs in the SSA22 field is consistent with previous studies, where median masses typically range from – M⊙ (cunha2015; Zavala et al., 2018; Dudzevičiūtė et al., 2020b; Shim et al., 2022; Hyun et al., 2023). Although the values are similar in magnitude, the median SFR in our SSA22 sample is 0.1–0.2 dex lower than in the literature.
From the AS2UDS sample (Dudzevičiūtė et al., 2020b), we selected sources at with radio or 24 counterparts. Their median stellar mass and star formation rate (SFR) are M⊙ and M⊙ yr-1, respectively. Applying the same redshift cut to the multi-band identified S2TDF sample (Hyun et al., 2023) yields median values of M⊙ for stellar mass and M⊙ yr-1 for SFR. Under identical criteria, we selected a conservative sample at with identifications from VLA or MIPS 24 (Table 5). This sample has a median stellar mass of M⊙ and a median SFR of M⊙ yr-1. While the SFR is consistent with the literature values cited above, the stellar mass is somewhat larger, a discrepancy that can likely be attributed to differences in SED fitting methodologies and results (see Section V). Our radio or 24 identified sample also includes a number of “quenched” systems, yet their SFRs are consistent with those reported in literature catalogs. If we further restrict the selection to sources at with radio or 24 counterparts that lie on or above the star-forming main sequence, the median stellar mass and SFR become M⊙ and M⊙ yr-1, respectively. Under the same conditions, the corresponding median values for the AS2UDS subsample are M⊙ and M⊙ yr-1, and for the S2TDF subsample are M⊙ and M⊙ yr-1. Therefore, we suggest that this may arise from relatively weak constraints in the far-infrared during SED fitting, which could lead to greater uncertainty in derived properties and consequently their misclassification.
The ADF22+ survey (Huang et al., 2025) has identified a population of relatively faint dusty star-forming galaxies (DSFGs) in the core of the protocluster. Within the redshift range , these galaxies have median stellar masses and SFRs of and , respectively—significantly lower than the corresponding medians for our SMGs in the same redshift interval, which are and . Nevertheless, their stellar mass and SFR volume density are nearly twice ours (Section V.1.1). This implies that conventional, shallower submillimeter/millimeter surveys miss a substantial population of faint dusty galaxies, yet these galaxies collectively contribute more to the star formation activity than the bright SMGs alone. Moreover, these DSFGs all lie on the star-forming main sequence, indicating that the majority of protocluster members are undergoing steady, gas-fueled star formation sustained by inflows from the cosmic web—rather than being driven by major mergers (Huang et al., 2025).
The median SFR and for the ‘MAIN’ and ‘SUPP’ samples are 178 M⊙ yr-1, M⊙ and 116 M⊙ yr-1, M⊙, respectively. However, we emphasize that the physical properties of sources identified through different methods show no significant differences. The median values of SFR, , , and are in close agreement, spanning ranges of 140-190 M⊙ yr-1, M⊙, L⊙, M⊙, respectively.
The median deblended 850 fluxes for counterparts of newly detected sources ( 2.8-3.5 mJy) and S2CLS-detected sources ( 3.5–4.2 mJy) are 3.2 mJy and 4.9 mJy, respectively. These two SMG groups show median SFR and values of 117 M⊙ yr-1, 1.59 1011 M⊙ and 186 M⊙ yr-1, 1.55 1011 M⊙, correspondingly. These results reveal systematic differences in SFR distributions across flux bins, the relation was derived by least-squares fitting: log10(SFR) = (0.49 0.21) log10(S) + (1.78 0.13). When restricting the sample to sources brighter than 4 (or 5) mJy, the median SFR increases to 192 (296) M⊙ yr-1. However, lower SNR sources do not significantly impact the estimated physical properties. For instance, sources with SNR exhibit median SFR and of 175 M⊙ yr-1 and M⊙, while lower-SNR sources (3.5-4) show values of 130 M⊙ yr-1, 1.39 1011 M⊙.
We adopt the main sequence relation from Speagle et al. (2014). We classify the SMGs exhibiting star formation rates three times above this relation as starburst systems, while those with values three times below are categorized as quenched systems (see Paper IV of this series). SMGs located within the main sequence have median SFR and values of 176 M⊙ yr-1 and M⊙, respectively. The ‘starburst system’ exhibit higher SFR but lower galaxy masses, with median values of SFR 510 M⊙ yr-1 and M⊙. In contrast, the ‘quench system’ are on average more massive, with a median of M⊙ — approximately twice the average of full sample — but a median SFR of only 28 M⊙ yr-1.
Overall, the SFR of SMGs evolves strongly with redshift, decreasing by nearly an order of magnitude over a 1 Gyr period when the universe was 2–3 Gyr old. Though this behaviour of increasing SFR () with redshift is very likely influenced by a well-know selection effect (miettinen2017; Dudzevičiūtė et al., 2020b). Fitting sources within , we obtain . In contrast, the stellar mass distribution remains relatively stable over a broad redshift range (), with a median mass of M⊙. At redshift above , SMGs exhibit median SFR M⊙ yr-1 and stellar mass M⊙. In the redshift range 1.5-3, median SFR decrease to approximately M⊙ yr-1 and stellar mass remains M⊙, respectively.
Following cosmic noon, massive galaxies gradually exhaust their gas reservoirs through duty cycles of intense star formation, transitioning out of the ‘SMG phase’ toward quiescent evolution. As a result, the number density of massive SMGs declines sharply at 2, rendering them increasingly difficult to detect. Concurrently, a notable reversal emerges at 1–2 (see Figure 19): lower-mass SMGs exhibit elevated SFRs and begin to dominate cosmic star formation activity. This intriguing phenomenon, known as the “downsizing” effect (Cowie et al., 1996), where star formation shifts from massive galaxies at earlier epochs to lower-mass systems at later times. Further discussion of these trends will be presented in the subsequent paper of this series (see Paper IV).
Below , 850 SMGs become increasingly challenging to detect, with the observed number density decline rapidly. Their median SFR and decline to 10 M⊙ yr-1 and M⊙, respectively. Cosmic star formation activity progressively shifts to other galaxy populations.
V.2.1 Mass-to-light Ratio
At typical redshifts for SMGs, the IRAC 5.8 band corresponds approximately to the rest-frame H-band, which serves as a reliable tracer of the dominant stellar population—low-mass stars (Dudzevičiūtė et al., 2020b). As shown in Figure 20, the 151 SMGs with 5.8 detections follow a distribution and trend consistent with the AS2UDS sample: brighter 5.8 emission corresponds to higher stellar mass. A linear fit yields: [M [mag] + (14.97 1.16).
The mass-to-light ratio characterizes the relationship between stellar mass and luminosity. We derive the rest-frame 1.6 luminosity from the best-fit SED for each source and estimate the H-band mass-to-light ratio, . Our sample exhibits a median of M⊙/L⊙, consistent with values from ALESS and AS2UDS.
cunha2015 suggest that mass-to-light ratio correlates with stellar mass, as the most massive galaxies tend to host older stellar populations and/or experience higher dust attenuation, resulting in larger mass-to-light ratio values. This implies a connection between galaxy mass and evolutionary state—more massive galaxies generally exhibit larger mass-to-light ratios. For instance, SMGs on the star-forming main sequence and those in quiescent phases show similar distributions, with relatively high stellar masses and elevated mass-to-light ratios, having median values of 1.61 and 3.83 M⊙/L⊙, respectively. In contrast, SMGs undergoing intense starburst activity—dominated by recently formed, young stars—exhibit significantly lower mass-to-light ratios, with a median 0.75 M⊙/L⊙,
Furthermore, shows strong correlations with both mass-weighted age (age) and dust attenuation () (Figure 20). The linear relations are: log , log .
Our median mass-weighted age, Myr, is consistent with that reported in AS2UDS. Older galaxies have a greater proportion of their light emitted by older stars, leading to a higher due to the lower luminosity efficiency of evolved stars. This trend underscores that not only the presence of low-mass stars, but also increased stellar evolutionary age, reduces luminous efficiency and raises the mass-to-light ratio.
Although the H-band is less affected by dust attenuation, dust obscuration still significantly influences the apparent . Among SMGs with mag (the sample average), the median is M⊙/L⊙, compared to M⊙/L⊙ for sources with lower extinction.
V.3 Infrared Luminosity
The median infrared luminosity (, integrated over 8–1000 in the rest frame) of the sample is L⊙, with a 1 range of – L⊙, and the most luminous source reaches L⊙ (lower-left panel of Figure 19). While slightly lower than literature values, the infrared luminosities of SSA22 SMGs remain comparable in order of magnitude to previous studies.
Similar to the SFR results discussed earlier, when we conservatively restrict our sample to sources with 1 that have radio or 24 counterparts, the median is 3.24 L⊙. For comparison, selecting AS2UDS sources that have multiwavelength counterparts and satisfy 1 yields a median infrared luminosity of 2.88 L⊙. Applying the same redshift cut to the S2TDF sample gives a median value of 3.63 L⊙, in excellent agreement with our result.
The majority of the SSA22 sample consists of ultra-luminous infrared galaxies (ULIRGs, L⊙) or hyper-luminous infrared galaxies (HLIRGs, L⊙), with 63% ( 8%) classified as ULIRGs (HLIRGs). At , SMGs predominantly fall into ULIRG category, while the rare, exceptionally bright HLIRGs primarily appear before , hosting some of the most intense star formation episodes in the universe. The remaining quarter of sources are luminous infrared galaxies (LIRGs, L⊙). These become increasingly prevalent at and represent nearly all SMGs below .
Due to the tight correlation between infrared luminosity and star formation activity, exhibits a significant evolution with the cosmic star formation history in Figure 19. Dudzevičiūtė et al. (2020b) found that over the redshift range –4, scales approximately as , but they argue this trend may be biased by the dependence of SED fitting on FIR data. Nevertheless, it is evident that the most luminous SMGs show a clear decline in brightness: HLIRGs are predominantly found at , and at , SMG infrared luminosities typically fall below L⊙.
is excellent indicator of star formation activity, and always shows a tight correlation with SFR, follows the relation (scaled to Chabrier (2003) IMF): SFR [M⊙ yr-1] 1.09 10-10 [L⊙] (kennicutt1998) , as shown by the cyan solid line in Figure 21. Our sample exhibits a scatter of approximately 0.3 dex around this relation.
We note several outliers with higher luminosities than predicted. At SFR M⊙ yr-1, deviations from this relation occur in systems with elevated AGN fractions (, derived from CIGALE fitting), with AGN luminosities typically 0.5–3 times, and up to 9 times, the dust luminosity. Coloring the data by mass-weighted age (age), we find that these AGN-dominated systems tend to be younger systems.
However, we note that sources residing in the large-scale structure, and those with significant AGN contributions do not systematically deviate from the infrared–SFR relation. We mark two confirmed X-ray AGNs (AGN7 and the quasar ADF22.A1; Lehmer et al., 2009b; Alexander et al., 2016; Monson et al., 2023; Umehata et al., 2025b), yet both lie well within the linear range of the relation between infrared luminosity and SFR. This is consistent with the findings of Monson et al. (2023), who concluded that the majority of AGN-hosting SMGs are indistinguishable from normal star-forming galaxies.
We also identify a group of counterparts with SFRs around 10 M⊙ yr-1 that lie significantly below the relation. Most sources at 1 do not show strong deviations; instead, the outliers are predominantly quiescent identifications. These typically exhibit mass-weighted stellar ages of 1–1.5 Gyr and a high fractional contribution from old stellar populations to the infrared luminosity. As noted by cunha2015; Dudzevičiūtė et al. (2020b), MAGPHYS allows heating of dust by older stellar populations, which can produce excess infrared emission—particularly in galaxies with high mass-weighted ages. This effect may contribute to the apparent offset of these quiescent systems from the canonical IR–SFR relation.
Finally, we find that sources with multiple counterparts do not systematically lie outside the expected relation. Given the low multiplicity fraction in our sample, we conclude that flux misallocation due to source blending does not significantly bias our results.
Therefore, we believe that both AGN activity and older stellar populations can cause galaxies to deviate from the canonical –SFR relation. AGN-dominated SMGs tend to be younger and more luminous, whereas systems with significant old stellar populations are generally older and located around . If these outliers are excluded, the median and SFR of the remaining sample are L⊙ and 253 M⊙ yr-1, respectively.
V.4 Dust and Dust fraction
Dust in galaxies originates from various processes including supernova ejection of massive star, stellar wind from evolved star, and infalling of intergalactic medium (IGM). While the dust is destructed through supernova shocks, feedback from stellar and AGN (liao2024). Dust grains efficiently absorb ultraviolet to near-infrared radiation and re-emit this energy in the infrared, it serves as a crucial tracer of star formation activity in DSFGs and a key constraint for galaxy evolution models. The dust-to-stellar mass fraction, , reflects the balance between dust production and destruction, and thus provides important insights into feedback processes in galaxies (Dudzevičiūtė et al., 2020a; liao2024).
The median dust mass () for SSA22 SMGs is M⊙, with the 16th–84th percentile range of – M⊙. Over 73% of the sources contain at least M⊙, with the most extreme reaching M⊙.
The dust mass estimates for our sample appear systematically higher than those reported in previous studies (cunha2015; Dudzevičiūtė et al., 2020b; Hyun et al., 2023). We attribute these differences to the distinct dust models employed: MAGPHYS utilizes modified blackbody spectra to model warm and cold dust components, whereas our CIGALE implementation adopts the dl2014 model, which accounts for dust heating by both star-forming regions and the diffuse interstellar radiation field. This interpretation is supported by liao2024, who report that CIGALE yields dust masses 1.3–1.5 times greater than MAGPHYS estimates. Furthermore, the inclusion of shorter FIR bands in SED fitting introduces systematic biases. Wavelength coverage near the far-infrared emission peak preferentially samples warmer dust components, thereby reducing sensitivity to colder dust and potentially underestimating total dust mass (tacconi2018). Although this effect will be less at high redshifts (tacconi2018), studies suggest a potential underestimate by a factor of 1.2–2 (birkin2021). Additionally, contamination during the identification process and potential misallocation of flux may also introduce uncertainties (Section III.3). While methodological variations exist in both approach and results, they do not materially impact subsequent analyses.
In general, at fixed infrared luminosity, SMGs with higher 850 flux densities tend to exhibit greater dust masses and lower temperatures (swinbank2014; cunha2015). As noted by Dudzevičiūtė et al. (2020b); Hyun et al. (2023), the 850 flux is tightly correlated with cold dust mass. We derive the following empirical relation for our sample: log10( []) = log10(S850 [mJy]) + . However, despite the correlation between and , and between and , no significant correlation is found between and .
Sources with log10(/M⊙) predominantly have log10(SFR/M⊙ yr-1) , while those with log10(SFR/M⊙ yr-1) generally exhibit log10(/M⊙) 9.2, suggesting an apparent anti-correlation between and SFR. However, given the strong evolution of SFR with redshift, this trend likely reflects the cumulative buildup of dust mass over time rather than a direct physical dependence. Overall, (and ) shows only a weak correlation with SFR (thus with ), and no clear relation with V-band attenuation (). This suggests that dust geometry and heating efficiency—rather than total dust mass—are more critical in determining the observed infrared luminosity of SMGs, as compact dust configurations yield higher temperatures and luminosities (cunha2015). Additionally, the diversity of SED and associated uncertainties in dust mass estimates further weaken any significant correlation between (or ) and /SFR.
V.4.1 Dust fraction
We define the dust mass fraction as . The median for the sample is , with a 16th–84th percentile range of –.
Figure 22 left panel illustrates the redshift evolution in : median values are at , at , and at . A linear fit to the full sample yields log . The slope is steeper than that reported by miettinen2017 for ALMA-detected SMGs in the COSMOS field. We believe this discrepancy arise from the incompleteness of high-redshift sources in our sample. When restricting our analysis to , the trend becomes consistent with previous results: decreases at and remains approximately flat at (miettinen2017; liao2024).
The dust mass fraction correlates with the evolutionary state of galaxies. While main-sequence SMGs follow a trend consistent with the overall sample, starburst and quenched systems deviate by approximately 0.5 dex. We note that SMGs in different evolutionary states possess distinct stellar mass distributions, indicating that the apparent correlation between and evolutionary phase is partially driven by its underlying relationship with stellar mass.
A notable transition in occurs at –2.5 for main-sequence SMGs (Figure 22 (b)), in good agreement with previous studies. Concurrently, the average of quenched systems shows an opposite trend, increasing across the same redshift range. We interpret this behavior as a signature of “downsizing”: a population of intermediate-mass, dust-rich main-sequence galaxies gradually transitions to quiescence at , causing a decline in the dust content of the main-sequence sample while increasing the average of the quiescent population. Over time, lower-mass starburst galaxies transition onto the main sequence, leading to a recovery in the dust fraction of the main-sequence population. This evolution proceeds dynamically, by , most intermediate-mass galaxies in the 850 field have become quiescent, and their dust content rises to levels comparable to those of main-sequence galaxies. Figure 22 panel (c)) shows that the values for starburst SMGs in SSA22 are higher than those reported by miettinen2017, but consistent with liao2024, particularly when accounting for differences in the definition of starburst galaxies.
We find strong correlations between and both and (Figure 23): log10, log10. Dust fractions increase with decreasing stellar mass, with median of 0.005, 0.014, and 0.047 for , , and M⊙, respectively. Moreover, at fixed stellar mass, galaxies at lower redshifts consistently exhibit higher dust fractions, further supporting the picture of dust content evolving with cosmic time.
We suggest that the observed trend of decreasing with redshift is jointly driven by the increase in dust mass and the decrease in stellar mass at lower redshifts. The decreasing evolve of stellar mass arise from the “downsizing” scenario. The increasing dust is deeply influenced by rising gas-phase metallicity (bethermin2015; liao2024), and shifts in the balance between metal enrichment, grain growth, and dust destruction (miettinen2017; Dudzevičiūtė et al., 2020a; liao2024). The gas-phase metallicity increases toward lower redshift (bethermin2015), implying that () and dust formation efficiency also rise in the same direction (miettinen2017; liao2024). Additionally, the rate of dust accumulation regulated by cold gas surface density, while AGN feedback potentially suppressing dust growth by reducing gas density (liao2024).
V.5 Dust attenuation
In the rest frame, ultraviolet to near-infrared photons from galaxies are attenuated by dust in both birth clouds and the diffuse interstellar medium. However, the estimated represents an average attenuation across the entire galaxy and does not account for the spatial distribution of stars and dust, and the actual attenuation become increasing severe at shorter wavelengths. Thus, should be regarded as a lower limit to the true attenuation (Dudzevičiūtė et al., 2020b). Nevertheless, remains a useful indicator for assessing the level of dust obscuration in these highly active DSFGs.
We measure the median for the sample to be mag, with 68% of sources lying within the range 1.69–4.44 mag. On average, this corresponds to a factor of 20 attenuation in the optical and near-infrared bands. Combined with the -correction at high redshift, this implies that constructing a mass-complete sample based solely on UV/optical observations is highly challenging, as heavily obscured galaxies are likely to be missed (wang2019; Dudzevičiūtė et al., 2020b).
Our median and its distribution are consistent with those reported for AS2UDS ( mag, Dudzevičiūtė et al., 2020b) and S2TDF ( mag, Hyun et al., 2023), and slightly higher than SCUBADive ( mag, McKinney et al., 2025), although the overall distributions remain comparable. Our value is also higher than that of ALESS( 1.9 mag, cunha2015), which may be due, as suggested by Dudzevičiūtė et al. (2020b), to cosmic variance leading to an underrepresentation of highly obscured SMGs at in their sample.
The median values for all VLA counterparts, 24 counterparts, 8 counterparts, and color-selected counterparts are 3.10, 2.65, 3.06, and 3.13 mag, respectively. Since most 8 and color-selected counterparts are also identified in the VLA or 24 samples—and radio-detected sources dominate the overall counterpart population—their dust attenuation remains consistent with that of the radio-identified sample. Color-selected galaxies do indeed exhibit redder colors and higher dust attenuation. Considering only those color-selected galaxies without counterparts from other identification methods, the median rises to 3.82 mag. However, this result could be influenced by contamination from lower-redshift sources. When restricted to , the median values for the respective counterpart samples become 3.10, 3.04, 3.08, and 3.09 mag.
Dividing the sample into optically bright and faint sources (based on whether the number of detected optical/NIR bands exceeds four), we find the median values are mag (bright) and mag (faint), respectively. As expected, optically faint sources experience significantly stronger dust obscuration, attenuating V-band light by a factor of on average—about three times more than bright sources.
The distribution of with redshift (Figure 24) shows no significant evolutionary trend. We confirm that the ‘MAIN’ and ‘SUPP’ subsamples exhibit statistically indistinguishable distributions, with no significant differences between them (Table 5). We note that SMGs’ attenuation decrease significantly at , the median is , consistent with other SMG surveys (e.g., AS2UDS, S2TDF), which also show a decline in with increasing redshift—reaching at .
Recent studies have analyzed high-redshift optically faint/dark galaxies (OFGs), which are defined as H-band dropouts and comparable optically faint populations (Williams et al., 2024). These investigations reveal that the majority of OFGs are extended dusty star-forming systems, with only a minor fraction being quiescent galaxies (Barrufet et al., 2025). Their SFRs typically range from 10 to 100 M⊙ yr-1, while stellar masses are approximately 1-2 dex lower than those of SMGs (Gottumukkala et al., 2024; Williams et al., 2024). These findings highlight OFGs as a critical population bridging extreme DSFGs and the broader population of massive galaxies, warranting further investigation.
We compare the sample with those OFGs. Although OFGs generally have lower SFRs, stellar masses, and compared to SMGs, they are still more dust-obscured, more actively star-forming, and more massive than typical optically selected galaxies. The OFG sample of CEERS and GOODS-S field (orange triangles in Figure 24; Gottumukkala et al., 2024; Williams et al., 2024) has an average of 2.7 mag, rising to 3.1 mag at , but decreasing significantly at to a median of 1.9 mag. This behavior closely resembles the trend seen in SMGs. Interestingly, JWST-selected optical galaxies888https://s3.amazonaws.com/aurelien-sepp/primer-cosmos-east-grizli-v7.0/catalog/primer-cosmos-east-grizli-v7.0_morpho-phot.fits.gz (transparent hexagons in Figure 24) at –5 exhibit a similar trend.
Not only SMGs, but also OFGs and UV-selected galaxies exhibit a similar trend. If this observed trend is not an artifact of sample selection or small-number statistics, we interpret it as evidence for a fundamental shift in the dominant mode of cosmic star formation—from an unobscured to a dust-obscured regime (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020; zavala2021). The underlying physical driver may be the gradual buildup of metallicity in galaxies over cosmic time, which eventually reaches a threshold enabling efficient dust production and obscuration (see Section V.4).
simpson2014; Dudzevičiūtė et al. (2020b) found that OFGs in the ALESS and AS2UDS samples tend to lie at higher redshifts. Smail et al. (2021) noted that, despite having dust masses and stellar masses comparable to those of other SMGs, these near-infrared–faint SMGs exhibit more compact dust emission and higher star formation rate surface densities. This suggests that such systems may be linked to the formation of compact, spheroidal structures observed in high-redshift galaxies.
The GOODS-S field OFGs sample (Xiao et al., 2023) has a median redshift of and relatively low average ( mag). Among these, sources detected by ALMA (red star Pattern in Figure 24) exhibit higher , dust mass, and SFR, along with more abundant and colder dust, suggesting that more heavily obscured systems tend to have colder dust temperatures. In contrast, Manning et al. (2025) identify a population of MIR-faint 1.5 -dropouts (i.e., F444W faint F150W-dropouts) in the COSMOS field (blue triangles in Figure 24), characterized by high redshifts and substantial dust attenuation, with median values of and mag, respectively. Most of these sources lie on the star-forming main sequence. However, due to the lack of FIR and (sub)millimeter constraints, estimated physical properties for such sources are subject to non-negligible uncertainties (Williams et al., 2024; McKinney et al., 2025).
The brightest 850 SMGs ( mJy) have an average mag, which seem suggesting a link between high submillimeter flux and extreme dust obscuration. However, this may reflect small-number bias, as no significant correlation is found between and across the full sample. In the large AS2UDS sample, the brightest SMGs show mag but remain consistent with the median . Galaxies in starburst or quiescent phases exhibit values close to the sample average, while main-sequence SMGs show slightly lower attenuation ( mag). Although more massive SMGs (log10(/M⊙) ) tend to have higher (typically mag), and those with have a slightly lower median ( mag), we find no significant correlation between and stellar mass, SFR, dust mass, or dust fraction.
The lack of correlation with key physical properties or star formation state suggests that dust geometry—rather than intrinsic galaxy properties—dominates the observed attenuation.
V.5.1 Attenuation and Optical Color
In this section, we briefly examine the relationship between and optical colors in SMGs. Given their substantial dust attenuation, we focus on the correlation between () color and .
Among the 104 sources in our sample with detections in both and bands, the () color ranges from 0 to 2 mag. Although the UDS deep field enables AS2UDS to probe significantly redder sources—up to () 3–4 mag—our () distribution is in good agreement with AS2UDS. The median redshift and for the -detected SMGs in SSA22 are 1.53 and 2.37 mag, respectively, while the remaining sources, as expected, exhibit higher median redshift and (2.16 and 3.86 mag). These sources undetected in the or bands likely occupy the regime of higher and redder () colors in the color–magnitude space.
At the median redshift of our sample (), the - and -bands correspond approximately to the rest-frame - and -bands, respectively. The attenuation in the rest-frame - and -bands, and can be expressed as:
From the color-excess relation:
we derive the linear relation between observed color and AV:
where (U-R) is the observed color, (U-R)0 is the intrinsic color, , describes the reddening law.
For redshifts 1–4, we compute the rest-frame wavelengths corresponding to and bands and estimate the intrinsic and using an average stellar spectrum. As highly star-forming galaxies, SMGs have strong intrinsic optical emission, resulting in modest intrinsic colors: , with a median of 0.78. Assuming the Calzetti extinction law (calzetti2000), ranges from 0.5 to 0.7, with a value of 0.60 at 2.
However, as shown in Figure 25, the observed values for SMGs are systematically higher than those predicted by the Calzetti law, indicating that this low-redshift prescription does not accurately describe the dust attenuation in high-redshift SMGs. The Calzetti reddening law is derived from low-redshift UV-bright star-forming galaxies, which differ fundamentally from the dust-rich SMGs. Furthermore, Menéndez-Delmestre et al. (2009) note that SMGs exhibit more extended star formation and dust distributions compared to local starbursts and ULIRGs, leading to different extinction properties.
Additionally, SMGs are typically selected based on flux in specific bands, resulting in a heterogeneous population (hayward2011; Johnson et al., 2013). For instance, galaxies selected in this way span over two orders of magnitude in both stellar mass and SFR, exhibit a wide range of dust masses and spatial distributions, and encompass diverse evolutionary states. The sample commonly includes both merging or interacting systems and isolated galaxies, with some hosting AGN while others do not. This diversity is evident in their SEDs and radio properties (see Paper IV for the discussions of SED and IR-radio correlation). Attenuation curves can vary significantly between different galaxy types, and even within a single galaxy, they may differ substantially along different lines of sight (uematsu2024). This intrinsic diversity further limits the applicability of a single, universal extinction law to the SMG population.
As noted earlier, the intrinsic rest-frame colors of SMGs are relatively modest; fixing this value at the median of 0.78 and fitting the data yields a slope of , corresponding to (for = 4.05; calzetti2000). This implies that highly obscured high-redshift star-forming galaxies require significantly steeper attenuation in the blue optical regime than predicted by the Calzetti law. A steeper reddening curve—relative to the canonical Calzetti form—is necessary to reproduce the observed dust attenuation in SMGs. This enhanced attenuation may be attributed to the fact that newly formed stars in SMGs are still embedded within dense molecular clouds, leading to more severe extinction of their strong UV emission, consistent with the two-component extinction model employed in our SED fitting (Section IV.2). However, the substantial scatter in the data also highlights the diverse SED properties of SMGs and/or variations in dust geometry, which may collectively produce a wide range of dust attenuation patterns.
VI Summary
In this work, we present a multiwavelength analysis of 850 SMGs in the SSA22 deep field.
Using four identification method (Figure 8), we associate 248 SMGs ( with a deblended 850 flux density of mJy) with 192 SCUBA-2 sources, of which 136 counterparts to 101 SCUBA-2 sources are reported here for the first time (Section III). We confirm the effectiveness of identification techniques based on IRAC 8 -band and optical/NIR color selection (Figure 12). The average multiplicity of SCUBA-2 sources is 26%, with brighter sources exhibiting higher multiplicity (Figure 13).
We apply the “super-deblending” technique (Section II.5, Appendix B) to obtain deblended FIR and submillimeter fluxes, and combine these with panchromatic photometry from optical to radio wavelengths to perform SED fitting using CIGALE (Section IV.2). After applying quality cuts based on SED fitting reliability, our final sample comprises 221 SMGs ( with a deblended S850 of mJy) associated with 186 SCUBA-2 sources. Our main results are summarized as follows:
1.
The median redshift of SSA22 SMGs is , with consistent distributions across identification methods (Figure 17). This is slightly lower than literature values, likely due to incompleteness in our multiwavelength data at the highest redshifts. The comoving volume density of SMGs increases by a factor of at , plateauing at 1.78-3.16 10-5 cMpc-3 over . Newly identified SMGs show a similar redshift distribution (median 1.9) to the overall sample. Approximately two-thirds of sources are classified as optically bright (median redshift ), while optically faint sources exhibit higher median redshift (), with completely undetected optical/NIR sources reaching . Brighter 850 SMGs are preferentially found at higher redshifts.
2.
SMGs in large-scale structures show significant overdensity (Figure 17 image (c)), demonstrating their reliability as traces of cosmic structure formation at high redshift. Additionally, the SMG overdensity exceeds blank field average by more than an order of magnitude (Section V.1.1; Figure 17 image (e)).
3.
The median stellar mass and SFR of our SMG sample are M⊙ and M⊙ yr-1, with 16-84% ranges of M⊙ and M⊙ yr-1, respectively (Figure 19). Newly detected SMGs have comparable stellar masses but lower SFRs, consistent with their fainter submillimeter fluxes. Starburst-phase SMGs exhibit high SFRs and relatively lower stellar masses, whereas quenched systems are more massive but have suppressed star formation.
4.
The stellar mass distribution of SMGs remains remarkably stable across cosmic time, while SFR evolution remains debated due to selection effects (miettinen2017; Dudzevičiūtė et al., 2020b). We observe a clear “downsizing” signature: at early epochs ( 2), both high- and low-mass SMGs undergo intense star formation. After cosmic noon (), massive SMGs exhaust their gas reservoirs and transition to quiescence, while lower-mass SMGs continue forming stars and dominate the cosmic SFR density.
5.
Older galaxies exhibit higher mass-to-light ratios due to increased contributions from evolved stellar populations and reduced luminosity efficiency (Figure 20). Dust attenuation and the presence of low-mass stars also significantly affect . Quiescent SMGs show systematically higher ratios than star-forming ones.
6.
The median infrared luminosity is (2.25 0.25) 1012 L⊙, with 68% range of 0.46-6.93 1012 L⊙ (Figure 19). Approximately 63% ( 8%) of the sample are classified as ULIRGs (HLIRGs). A clear decline in brightness is observed among the most luminous SMGs, indicating significant evolution of infrared luminosity with cosmic time. At , nearly all SMGs are ULIRGs, while HLIRGs primarily appear before . Although is tightly correlated with SFR, deviations from this relation can occur: AGN activity in young SMGs and excess emission from evolved stellar populations in older systems can drive significantly above the canonical relation (Figure 21).
7.
The median dust mass is (1.95 0.14) 109 M⊙ (16-84% range of 0.54-5.31 109 M⊙), slightly higher than literature values–possibly due to differences in dust modeling assumptions. The 850 flux strongly traces cold dust mass, whereas is governed more by dust geometry, compactness, and heating efficiency than by total dust mass or dust fraction.
8.
The median dust fraction is (1.4 0.2) 10-2, with 68% value range of . Dust fraction decrease at but remains roughly flat at (Figure 22). This evolution is shaped by the combined redshift trends of dust and stellar mass (Figure 23). Dust mass growth is regulated by rising gas-phase metallicity, cold gas surface density, and the balance between grain growth and destruction. Stellar mass evolution follows the “downsizing” scenario, which also drives a turnover in at , with decreases for main-sequence SMGs and increases for quiescent systems.
9.
The median is 3.09 0.07 mag (68% range: 1.69-4.44 mag). Optically bright and faint SMGs have mag and mag, respectively. While shows no strong evolution over , we detect a significant drop in attenuation at –not only for SMGs, but also for OFGs and UV-selected galaxies (Figure 24). This phenomenon possibly reflecting a transition from unobscured to dust-obscured star formation modes (Bourne et al., 2017; Dunlop et al., 2017; Bouwens et al., 2020; zavala2021), potentially driven by metallicity accumulation thresholds.
10.
shows no clear correlation with stellar mass or evolutionary state, suggesting that dust geometry–not just total dust content–dominates attenuation. Dust attenuation with observed () color, but the scatter and offset from predictions based on the Calzetti law indicate that a single, local reddening law is insufficient for high-redshift SMGs (Figure 25). Their diverse SEDs and composite nature further complicate attenuation modeling.
Acknowledgments
We sincerely thank Ian Smail for his constructive comments and insightful suggestions, which significantly improved the rigor and clarity of this manuscript.
Y.A. acknowledges the support from the National Key R&D Program of China (2023YFA1608204), the National Natural Science Foundation of China (NSFC grant 12173089), the China Manned Space Program with grant no. CMS-CSST-2025-A10, and the Strategic Priority Research Program of the Chinese Academy of Sciences (grant No.XDB0800301).
K.M. acknowledges the Waseda University Grant for Special Research Projects (Project number: 2025C-484).
The James Clerk Maxwell Telescope is operated by the East Asian Observatory on behalf of The National Astronomical Observatory of Japan; Academia Sinica Institute of Astronomy and Astrophysics; the Korea Astronomy and Space Science Institute; the National Astronomical Research Institute of Thailand; Center for Astronomical Mega-Science (as well as the National Key R&D Program of China with No. 2017YFA0402700). Additional funding support is provided by the Science and Technology Facilities Council of the United Kingdom and participating universities and organizations in the United Kingdom and Canada. Additional funds for the construction of SCUBA-2 were provided by the Canada Foundation for Innovation.
This work is based on observations taken with the JCMT under project codes M15AI91 and MJLSC02. The MJLSC02 data are part of the SCUBA-2 Cosmology Legacy Survey. This research has made use of the NASA/IPAC Infrared Science Archive, which is funded by the National Aeronautics and Space Administration and operated by the California Institute of Technology. This research has made use of the SVO Filter Profile Service “Carlos Rodrigo”, funded by MCIN/AEI/10.13039/501100011033/ through grant PID2023-146210NB-I00.
Appendix A Optical/Near-infrared Data Reduction
We use SExtractor (Bertin and Arnouts, 1996) to detect sources and perform photometry on the optical/near-infrared images. The full set of configuration parameters is listed in Table 6. This section primarily references Bertin and Arnouts (1996); Holwerda (2005) and the SExtractor official documentation 999https://sextractor.readthedocs.io/en/latest/.
Source detection and photometry are carried out on sky-background-subtracted images, following the background estimation procedure described in Bertin and Arnouts (1996); Holwerda (2005). SExtractor divides the image into a grid of cells with size defined by BACK_SIZE. Within each cell, it iteratively computes the median and standard deviation of pixel values, rejecting outliers beyond from the median until convergence. The final background value for each cell is determined based on local source density: in sparse regions—identified when the change in during iteration is less than 20%—the mean is adopted; otherwise, the mode is estimated as . A median filter of size BACK_FILTERSIZE is then applied to suppress residuals from bright or extended objects. Finally, a full-resolution background map is generated via spline interpolation and subtracted from the original image.
To improve photometric accuracy, we set BACKPHOTO_TYPE = LOCAL and specify the annulus thickness with BACKPHOTO_THICK. This allows SExtractor to estimate the local background residual from a rectangular ring surrounding each source and apply a correction to aperture photometry.
The RMS (background noise) map is constructed simultaneously from the per-cell standard deviations. We identify candidate sources as connected groups of at least 5 pixels exceeding above the local background (Hayashino et al., 2004; Umehata et al., 2014). The final source catalog is further refined by requiring that the peak pixel of each source exceeds . During detection, we employ the default conical convolution kernel to enhance the detectability of faint sources, although Holwerda (2005) suggests that a top-hat kernel may be more effective for very low-surface-brightness objects.
Each detected source is represented as a hierarchical tree of nested isophotal levels (typically 30, set by DEBLEND_NTHRESH). Deblending proceeds top-down: at each branching node with threshold , the object is split if (1) the flux in one branch exceeds a fraction (DEBLEND_MINCONT) of the total flux, and (2) at least one other branch also satisfies condition (1) (Bertin and Arnouts, 1996). Additionally, SExtractor performs source “cleaning” by evaluating whether neighboring objects would still be detected if the current source were removed—a critical step for recovering faint sources near bright stars or galaxies.
For photometry, we adopt MAG_AUTO as our fiducial flux estimator (Uchimoto et al., 2012). This method implements an adaptive Kron-like elliptical aperture that robustly captures the majority of a source’s flux across a wide dynamic range, even in crowded fields or in the presence of nearby contaminants. It effectively balances systematic and random photometric errors and provides a reliable estimate of the total magnitude (Bertin and Arnouts, 1996; Holwerda, 2005), making it the standard choice in extragalactic surveys.
The MAG_AUTO algorithm is based on the “first-moment” technique introduced by Kron (1980). An initial ellipse is defined from the second moments of the light distribution and scaled by a factor of 6 along both axes. Within this enlarged region, the Kron radius is computed as
A Kron factor is then applied to define the photometric aperture semi-axes as and , where is the ellipticity derived from the initial moment analysis. For sources with Gaussian-like profiles (e.g., stars and typical galaxies), Kron (1980) and Infante (1987) showed that encloses 90% of the total flux, nearly independent of brightness. Increasing to 2.5 raises this fraction to 94%, albeit at the cost of reduced signal-to-noise. We adopt to maximize flux recovery while accepting a modest decrease in photometric precision.
To prevent underestimation of flux for faint sources in high-noise regions, we enforce a minimum Kron radius. Furthermore, to mitigate contamination from overlapping or nearby bright/extended sources, we set MASK_TYPE = CORRECT, which replaces contaminated pixels within the aperture with symmetric, uncontaminated counterparts from the opposite side of the source center.
| Procedure | Parameter | Value |
| Background | BACK_TYPE | AUTO |
| BACK_SIZE | 64 | |
| BACK_FILTERSIZE | 3 | |
| BACKPHOTO_TYPE | LOCAL | |
| BACKPHOTO_THICK | 24 | |
| Detection | THRESH_TYPE | RELATIVE |
| DETECT_TYPE | CCD | |
| DETECT_MINAREA | 5 | |
| DETECT_THRESH | 2 | |
| FILTER | Y | |
| FILTER_NAME | default.conv | |
| DEBLEND_NTHRESH | 32 | |
| DEBLEND_MINCONT | 0.005 | |
| CLEAN | N | |
| CLEAN_PARAM | 1.0 | |
| MASK_TYPE | CORRECT | |
| Photometry | PIXEL_SCALE | 0.2 |
| PHOT_AUTOPARAMS | , | |
| PHOT_AUTOAPERS | , | |
| SATUR_LEVEL | 50000.0 | |
| SATUR_KEY | SATURATE |
Appendix B Super-deblending
The FIR information of galaxies is crucial for measuring the star formation rates and dust properties of galaxies (Liu et al., 2018; Zavala et al., 2018). However, FIR/mm images have coarse spatial resolution ranging from 14 - 35, leading to significant confusion, which not only blurs our observations of the deep universe but also hinders our understanding of the characteristics of high-redshift star-forming galaxies. To obtain deblended source fluxes from 250 to 1.1 mm, we adopted an innovative technique proposed by Liu et al. (2018); Jin et al. (2018), “super-deblended”, to deblend these images and obtain accurate FIR fluxes for SMGs. We will provide a brief introduction to the super-deblended process, while detailed principles and steps can be found in Liu et al. (2018); Jin et al. (2018).
The galaxy main sequence ensures that FIR/mm detected objects always have counterparts in IRAC images (Liu et al., 2018). Due to the high sensitivity of Spitzer/IRAC imaging and the availability of deeper IRAC advanced productions, we used MIPS 24 and radio sources with IRAC-detected counterparts to create a prior source catalog. We combined the SEIP and SpIES catalogs to obtain the IRAC source catalog and matched it with the VLA 3 GHz catalog. We selected 24 or radio sources with SNR greater than 3 as prior sources. Additionally, we utilized “the representative redshift catalog” in the SSA22 field compiled by Mawatari et al. (2023) and available photometric redshift information (see Section IV.1) to identify sources with low SNR ( 2.5) but located at higher redshifts ( 2) as part of the prior sources. The final source catalog includes 2683 “24+radio” prior sources.
The SED templates employed a combination of four component models, including stellar, mid-infrared AGN, dust continuum, and radio components. Additionally, the SED fitting also considering starburst populations. We introduce three flags to differentiate the fitting process for different types of galaxies: the MS/SB (main sequence or starburst galaxies) flag, the radio-excess flag, and the FIR photometry high-quality flag. The latter is used to prevent fitting failures when dealing with radio-loud or AGN sources. For detailed template information and the SED template fitting process, please refer to Liu et al. (2018).
The beam area of each FIR/mm image contains a significant number of star-forming galaxies, making it difficult to obtain an effective deblended flux when considering the contribution of all sources within the beam. Therefore, we set a threshold flux S, by considering the average density of prior sources, i.e. the number of prior sources within each beam. The threshold flux S is set within the range of 1 to 3, ensuring that the number of fitted sources in each band remains at a reasonable value of about or less than one per PSF beam area, see also Figure 26. If the sum of the prior source SED fitting flux and twice the fitting uncertainty, S + 2S, is less than S, we exclude the prior source in this band due to the faintness of the sources. We subtract these faint sources from the observed images and then use galfit (Peng et al., 2010) for photometry measurement of the prior sources in the subtracted images (Figure 27).
Before obtaining photometry of the prior sources in the current band using galfit, we need to consider that although the 24+radio catalog provides good FIR deblending priors, there are still sources that contribute to the FIR flux but are not included in the prior catalog due to different galaxy properties. Therefore, for the residual images after prior source extraction, we perform “blind-extraction” using SExtractor with a detection threshold set to 2.5-3.0. After visual inspection, these sources are added to the prior source catalog, and photometry measurement is redone using galfit for the FIR observed images. The deblending process proceeds from the shortest to the longest wavelength band, one band at a time. During the super-deblending process for each band, we flag the prior sources and use the deblended photometry measured in the previous band to optimize the prior sources in the current band, and re-perform FIR/mm SED fitting for each source. This is a crucial step in super-deblended photometry, helping us obtain robust deboosted fluxes for the prior sources.
Finally, we use Monte Carlo simulation to verify whether super-deblended photometry has significant bias and correct the flux based on the simulation results. Furthermore, we provide uncertainty estimates for deblended photometry using a highly optimized and nearly ideal uncertainty estimation method, which is a key advantage of super-deblending approach. In fact, in the presence of highly correlated noise and confusion noise in FIR/mm images, galfit cannot provide accurate photometry uncertainty. In the faint source-excluded image mentioned above, we inject simulated sources at random positions, simulating one source each time. Then, sources are extracted for the simulated image, and this simulation process is repeated approximately 3000 times. Thus, we obtain input and output fluxes of the simulated sources, and galfit photometric errors. Additionally, we measure the local rms noise value of the simulated sources, the local absolute flux density and local scatter in the residual images, and the crowdedness parameter. These simulated parameters are used to correct the bias in super-deblended photometry and the final flux uncertainty.